### v2

# DeepFRC: An End-to-End Deep Learning Model for Functional Registration and Classification

Abstract
Functional data - observations in the form of curves or trajectories - arise in diverse domains such as biomedical sensing, motion capture, and handwriting recognition. A core challenge in functional data analysis (FDA) is accounting for phase variability, where misaligned temporal patterns hinder accurate inference. We introduce DeepFRC, an end-to-end deep learning framework for joint functional registration and classification. Unlike conventional approaches that decouple alignment and prediction, DeepFRC integrates class-aware elastic warping and a learnable basis representation into a unified architecture. This design enables temporal alignment and dimensionality reduction to be jointly optimized with classification, improving both interpretability and accuracy. We establish the first theoretical connection between alignment quality and generalization error, and validate our model on synthetic and real-world benchmarks. DeepFRC consistently outperforms state-of-the-art methods, especially in scenarios with complex temporal misalignment. Code is available at: https://github.com/Drivergo-93589/DeepFRC.
11footnotetext: † These authors contributed equally.22footnotetext: * Corresponding author: zengpch@shanghaitech.edu.cn
Keywords: Functional Data Analysis; Deep Learning; Registration; Classification
1 Introduction
Functional data analysis (FDA) is a key area for analyzing data that varies continuously over domains like time, space, or other variables (Ramsay and Silverman, 2005; Ferraty and Vieu, 2006; Srivastava and Klassen, 2016). Functional data is ubiquitous in fields such as biomechanics, neuroscience, healthcare, and environmental science, appearing in datasets like growth curves, wearable device signals, EEGs, fMRI scans, and air pollution levels. Despite its wide applicability, FDA faces challenges stemming from the infinite-dimensional nature of functional data, as well as issues related to smoothness and misalignment, necessitating advanced analytical tools (Wang et al., 2016; Matuk et al., 2022).
1.1 Motivation
Two key tasks in FDA are functional registration (curve alignment) and classification. Registration aligns curves to remove phase variability, enabling meaningful comparisons (Srivastava et al., 2011), while classification assigns labels based on underlying curve features. Traditionally, these tasks are addressed separately, with pre-registration followed by classification. However, this decoupled approach is inefficient, as the label significantly influences the progression pace of the curves, and alignment can provide valuable insights for classification (Liu and Yang, 2009; Tang et al., 2022). Jointly analyzing curve alignment and classification accounts for both temporal and structural variations, offering a more comprehensive understanding of the data.
Deep learning has revolutionized data analysis across various domains by enabling automatic feature extraction, representation learning, and scalability (LeCun et al., 2015). In FDA, deep learning presents an opportunity to integrate registration and classification into a unified framework. However, its application remains underexplored, with existing studies primarily focused on improving either registration accuracy (Chen and Srivastava, 2021) or classification performance (Yao et al., 2021; Wang and Cao, 2024), but rarely addressing both simultaneously. To address this gap, we propose an end-to-end deep learning framework that combines functional data registration and classification, eliminating the need for separate preprocessing while leveraging neural networks to model the complex, non-linear relationships inherent in functional data.
1.2 Related Work
Functional registration is a critical step in FDA, aimed at aligning functional data to correct for phase variability (Ramsay and Silverman, 2005). Traditional approaches include landmark-based methods (Kneip and Gasser, 1992; Ramsay and Silverman, 2005), metric-based methods (Wang and Gasser, 1997; Ramsay and Li, 1998; Srivastava et al., 2011; Srivastava and Klassen, 2016), and model-based methods (Tang and Müller, 2008; Claeskens et al., 2010; Lu et al., 2017). While effective in some contexts, these methods face limitations such as high computational cost, manual intervention, and sensitivity to noise. Landmark-based methods, for example, require subjective and potentially impractical landmark selection (Marron et al., 2015). Recently, neural network-based methods (Chen and Srivastava, 2021) have been proposed, but they are still typically used as preprocessing steps, disconnected from the downstream tasks.
Functional classification has traditionally relied on statistical methods like generalized functional regression (James, 2002; Müller, 2005) and functional principal component analysis (fPCA) (Hall et al., 2000; Leng and Müller, 2006), which involve dimension reduction followed by classification. However, these methods are limited by their reliance on handcrafted features and basis function selection, which restricts their adaptability to complex, non-linear data. The integration of deep learning into functional classification, as seen in works by Thind et al. (2020), Yao et al. (2021), and Wang and Cao (2024), addresses these issues, although they typically treat registration as a separate preprocessing task.
Few studies have integrated registration and classification within a unified framework for FDA. Lohit et al. (2019) proposed a Temporal Transformer Network (TTN) for time series classification, but it is not designed for functional data, neglecting the curve’s smoothness and infinite-dimensional nature, thus struggling with accurate alignment. More recently, Tang et al. (2022) introduced a two-level model for joint registration and classification, modeling warping functions via a mixed-effects approach. However, this model is heavily reliant on assumptions, computationally expensive, and not easily extendable to multi-class settings.
1.3 Contributions
This work introduces a unified framework for functional registration and classification with the following key contributions: (1) Model Innovation: We propose (a) a learnable basis representation that preserves infinite-dimensional structure, (b) a class-aware registration loss tailored for multi-class alignment, and (c) a joint training strategy that enables mutual reinforcement between alignment and prediction - surpassing prior approaches (Lohit et al., 2019; Chen and Srivastava, 2021). (2) Theoretical Foundation: We establish the first theoretical connection between registration fidelity and classification generalization, offering new insights into how structured alignment impacts downstream learning. (3) Empirical Validation: Extensive experiments on both synthetic and real datasets demonstrate the model’s ability to jointly optimize registration and classification. DeepFRC consistently outperforms state-of-the-art baselines, with ablation studies validating the necessity of its integrated design. The paper follows the order of contributions outlined above.
2 The Model
Consider the data $\big{\{}(x_{i}(\bm{t}_{i}),y_{i})\big{\}}_{i=1}^{N}$, where $x_{i}(\bm{t}_{i})=\big{(}x_{i}(t_{i1}),\dots,x_{i}(t_{in})\big{)}^{\intercal}$ represents the observed values of functional data $x(t)$ at $n$ time points for the $i$-th sample, and $y_{i}\in\{1,\dots,C\}$ denotes its class label. The functional data $x(t)$ is assumed to be a mean-square continuous process defined over $t\in[0,1]$, and the observations are misaligned. We hypothesize the existence of an underlying nonlinear time warping map $\gamma:t\mapsto\tilde{t}$ and a prediction function $f:x(\tilde{t})\mapsto y$. Our goal is to estimate $\gamma$ and $f$ simultaneously by an end-to-end deep learning model. This allows us to algin new functional data $x^{*}(\bm{t}^{*})$ using $\gamma(\bm{t}^{*})$ and classify it with $y^{*}=f(x^{*}(\gamma(\bm{t}^{*})))$. In the following, we detail the development of our deep neural networks to approximate both mappings $\gamma$ and $f$.
2.1 1D CNN for Time Warping
Our approach firstly utilizes 1D convolutional neural networks (1D CNNs) to extract temporal features from functional data, capitalizing on their demonstrated effectiveness in capturing temporal patterns (Kiranyaz et al., 2016; Ince et al., 2016; Chen and Srivastava, 2021; Kiranyaz et al., 2021). These networks process observed functional data points to produce latent features, which are then used to construct time-warping functions, enabling better registration of functional data.
Given an input sequence $x_{i}(\bm{t}_{i})\in\mathbb{R}^{n}$, we extract time-warping features $\tau(x_{i}(\bm{t}_{i}))\in\mathbb{R}^{n}$ using a multi-layer 1D CNN. Each layer applies a 1D convolution followed by a ReLU activation:
$\bm{h}^{(l)}=\text{ReLU}(W^{(l)}\circledast\bm{h}^{(l-1)}+B^{(l)}),\quad l=1,\ldots,l_{1}-1,$ (1)
where $\bm{h}^{(0)}=x_{i}(\bm{t}_{i})$. The final layer is a fully connected transformation:
$\tau(x_{i}(\bm{t}_{i}))=\text{ReLU}(W^{(l_{1})}\bm{h}^{(l_{1}-1)}+\bm{b}^{(l_{1})}).$ (2)
Each convolutional layer has learnable weights $W^{(l)}$ and biases $B^{(l)}$. Following Chen and Srivastava (2021), we set $l_{1}=4$ and use three hidden 1D‐CNN blocks (kernel size $3$; channels $16\!\rightarrow\!32\!\rightarrow\!64$). This encoder, parameterised by $\Theta_{1}$, outputs the warping field $\tau(x_{i}(\bm{t}_{i}))$ that captures temporal structure.
To construct a boundary-preserving diffeomorphic warping function, the latent features $\tau(x_{i}(\bm{t}_{i}))$ are transformed as:
$\tilde{\gamma}_{i}(t_{ij})=\frac{\sum_{\mu=0}^{j}\tau^{2}_{i\mu}}{\sum_{\nu=0}^{n}\tau^{2}_{i\nu}},\hskip 5.69054ptj=1,\dots,n,$ (3)
where $t_{i0}=0$, $t_{in}=1,\tau_{i0}=0$, and $\tau_{i\mu}(\mu>0)$ is the $\mu$-th element of $\tau(x_{i}(\bm{t}_{i}))$. This enures that $\tilde{\gamma}_{i}(0)=0,\tilde{\gamma}_{i}(1)=1$ and $\frac{d(\tilde{\gamma}_{i}(t))}{dt}>0$, satisfying the requirements of a valid warping function (Chen and Srivastava, 2021). To reduce roughness and preserve the geometric structure, the warping function is further smoothed:
$\gamma_{i}(t_{ij})=\frac{\sum_{\mu=0}^{j}\tilde{\gamma}_{i\mu}}{\sum_{\nu=0}^{n}\tilde{\gamma}_{i\nu}}.$ (4)
This smoothing step ensures a smooth and well-structured warping function (Chen and Srivastava, 2021).
2.2 Basis Representation for Aligned Function
By applying the time warping $\gamma_{i}(t_{ij})$ to the functional data $x_{i}(\bm{t}_{i})$, the aligned functional data $\tilde{x}_{i}(t)$ can be obtained using 1D linear interpolation on the points $\big{\{}\big{(}\gamma_{i}(t_{ij}),x_{i}(t_{ij})\big{)}\big{\}}_{j=1}^{n}$:
$\begin{split}\tilde{x}_{i}(t)\triangleq\sum_{r=0}^{n-1}\Big{[}x(t_{ir})+\frac{x_{i}(t_{i(r+1)})-x_{i}(t_{ir})}{\gamma_{i}(t_{i(r+1)})-\gamma_{i}(t_{ir})}(t-\gamma_{i}(t_{ir}))\Big{]}\cdot\mathds{1}_{\{\gamma_{i}(t_{ir})\leq t\leq\gamma_{i}(t_{i(r+1)})\}},\end{split}$ (5)
where $\mathds{1}_{\{\cdot\}}$ is the indicator function. We choose this linear interpolation because it is both stable and computationally efficient (Berrut and Trefethen, 2004).
While traditional methods often discretize $\tilde{x}_{i}(t)$ into a high-dimensional vector $\tilde{\bm{x}}_{i}$ for input into neural networks, this approach suffers from the curse of dimensionality and fails to capture the smoothness of functional data, especially in noisy settings (Yao et al., 2021). To address these issues, we approximate $\tilde{x}_{i}(t)$ using basis expansion:
$\tilde{x}_{i}(t)\approx\sum_{j=1}^{K}c_{ij}\phi_{j}(t),$ (6)
where $\{\phi_{j}(t)\}_{j=1}^{K}$ are set to Fourier basis functions and $\{c_{ij}\}_{j=1}^{K}$ are coefficients to estimate. Using least squares, the coefficients $\bm{c}_{i}\triangleq(c_{i1},\dots,c_{iK})$ are estimated as:
[…]
2.3 Fully Connected Neural Network Classifier
To classify the $i$-th sample, we use a fully connected feedforward neural network that takes the input feature vector $\tilde{\bm{c}}_{i}\in\mathbb{R}^{K}$. The network has $l_{2}-1$ hidden layers with ReLU activations, followed by a softmax output layer:
$\bm{h}^{(l)}=\text{ReLU}(W^{(l)}\bm{h}^{(l-1)}+\bm{b}^{(l)}),\quad l=l_{1}+1,\dots,l_{1}+l_{2}-1.$ (8)
Here, $W^{(l)}$ and $\bm{b}^{(l)}$ are learnable weights and biases. The output $\psi(\tilde{\bm{c}}_{i})\in\mathbb{R}^{C}$ gives the class probabilities. The predicted label is:
$\psi(\tilde{\bm{c}}_{i})=\text{Softmax}(W^{(l_{1}+l_{2})}\bm{h}^{(l_{1}+l_{2}-1)}+\bm{b}^{(l_{1}+l_{2})}).$ (9)
The classifier is a three-layer MLP ($l_{2}=3$) with $8$ and $4$ hidden units, parameterised by $\Theta_{2}$.
2.4 Objective Function Formulation
We denote the overall parameters to estimate as $\Theta=\{\Theta_{1},\Theta_{2}\}$. Given training data $\{(x_{i}(\bm{t}_{i}),y_{i})\}_{i=1}^{N}$ (where $y_{i}\in\{1,\dots,C\}$), the objective function consists of two parts for estimating $\Theta$.
Class-aware Registration Loss.
We adopt the elastic Riemannian metric square-root velocity function (SRVF) (Srivastava and Klassen, 2016) for registration, which overcomes limitations of the $\mathds{L}^{2}$ metric such as lack of invariance to warping, asymmetry, and the pinching effect (Marron et al., 2015; Srivastava and Klassen, 2016). The SRVF of $x(t)$ is defined as $q(t)=\text{sign}(\dot{x}(t))\sqrt{|\dot{x}(t)|}$, and for the warped function $x(\gamma(t))$, it becomes $(q\star\gamma)(t)\triangleq q(\gamma(t))\sqrt{\dot{\gamma}(t)}$ (Joshi et al., 2007; Srivastava and Klassen, 2016). For the $i$-th sample, we denote the observed SRVF as $Q_{i}(\bm{\gamma}_{i})=\big{(}(q_{i}\star\gamma_{i})(t_{i1}),\dots,(q_{i}\star\gamma_{i})(t_{in})\big{)}^{\intercal}$. Inspired by the work of Chen and Srivastava (2021), we define the class-aware registration loss as:
$\begin{split}\mathcal{L}_{1}(\Theta_{1})={{\sum}}\limits_{j=1}^{C}\frac{\sum\limits_{i\in\{y_{i}=j\}}\lVert Q_{i}(\bm{\gamma}_{i})-\bar{Q}^{(j)}\rVert}{N^{(j)}}+\alpha\sum\limits_{1\leq u<v\leq C}\lVert\bar{Q}^{(u)}-\bar{Q}^{(v)}\rVert^{-1},\end{split}$ (10)
[…]
Classification Loss.
The cross-entropy loss is given by
$\mathcal{L}_{2}(\Theta)=-\frac{1}{N}\sum\limits_{i=1}^{N}\sum_{j=1}^{C}y_{ij}\text{log}\psi_{ij},$ (11)
where the label $y_{i}$ is represented as a one-hot encoded vector $(y_{i1},\dots,y_{iC})^{\intercal}$ where all elements are set to zero except for the element that corresponds to the active class, which is set to one.
The combined objective is:
$\mathcal{L}(\Theta)=\mathcal{L}_{1}(\Theta_{1})+\beta\mathcal{L}_{2}(\Theta),$ (12)
where $\beta$ balances registration and classification. Minimizing $\mathcal{L}_{1}$ aligns functions across classes while aiding classification, especially for similar classes, while minimizing $\mathcal{L}_{2}$ predicts class labels and improves alignment by enhancing class separation.
Integrating Subsections 2.1-2.4, we propose an end-to-end deep learning model, termed DeepFRC, for simultaneous functional registration and classification (Figure 1).
Figure 1: Overview of DeepFRC. The light blue region highlights the 1D convolutional module. The three blue arrows (left to right) denote: flattening of the $(l_{1}{-}1)$th layer matrix, the time warping (TW) operation, and the basis projection of the aligned function. “TW” and “DR” denote time warping and dimensionality reduction, respectively. We set $l_{1}=4,l_{2}=3$.
2.5 Optimization, Model Selection and Computational Complexity
The objective in Eq. (12) is optimized via stochastic gradient descent over the model parameters $\Theta$, using the AdamW optimizer (Ilya and Frank, 2019) for efficient and stable convergence, particularly in deep architectures (Diederik and Lei, 2015). Gradients with respect to each $\theta\in\Theta$ are computed via the chain rule, as detailed in Eqs. (A.1)–(A.2) of Appendix A. The complete training procedure is outlined in Algorithm 1.
The number of basis functions $K$ is empirically set to 100 and shown to be robust across tasks (Section 4.2). Hyperparameters $\alpha$ and $\beta$, which balance alignment and classification loss terms, are selected via data-splitting methods following (Wang et al., 2023); further model selection details are provided in Appendix B.
DeepFRC achieves linear time complexity $\mathcal{O}(Nn)$ with respect to sample size $N$ and sequence length $n$. In contrast, traditional alignment techniques such as dynamic time warping (DTW), which rely on dynamic programming, incur quadratic cost $\mathcal{O}(Nn^{2}k)$ (with $k<n$), rendering them computationally infeasible for long sequences (Sakoe and Chiba, 1978; Chen and Srivastava, 2021).
Algorithm 1 Training DeepFRC
0:  Training data $\big{\{}(x_{i}(\bm{t}_{i}),y_{i})\big{\}}_{i=1}^{N}$
1:  Set Hyperparameters: Size of basis function $K$, loss-related $\{\alpha,\beta\}$, and training parameters $\bm{\eta}$ (epochs $E$, batch size, learning rate, etc.)
2:  Initialize Parameters $\Theta=\Theta_{\text{initial}}$
3:  for $e=1$ to $E$ do
4:     Forward Propagation:(1) Compute $\gamma_{i}(\bm{t}_{i})$ for each $x_{i}(\bm{t}_{i})$(2) Warp $x_{i}(\bm{t}_{i})$ to obtain $\tilde{x}_{i}(t)$, calculate its SRVF $Q_{i}(\bm{\gamma}_{i})$, and extract coefficients $\tilde{\bm{c}}_{i}$(3) Pass $\tilde{\bm{c}}_{i}$ through the classifier to compute $\psi(\tilde{\bm{c}}_{i})$(4) Compute the loss $\mathcal{L}(\Theta)$
5:     Backward Propagation: Update $\Theta$ via AdamW optimizer using $\frac{\partial\mathcal{L}(\Theta)}{\partial\theta},\theta\in\Theta$
6:  end for
7:  Return Trained DeepFRC with optimized parameters $\Theta^{*}$
3 Theoretical Analysis
We now discuss the theoretical properties of the proposed model, demonstrating that it achieves low registration error under elastic functional data analysis (EFDA) (Srivastava and Klassen, 2016) framework and low generalization error under mild regularity conditions. Detailed proofs are provided in Appendices C and D.
The phase variability is modeled using monotone and smooth transformations within the set $\Gamma=\{\gamma:[0,1]\to[0,1]\mid\gamma(0)=0,\gamma(1)=1,\dot{\gamma}>0\}$. Under the framework of elastic functional data analysis (EFDA) (Srivastava et al., 2011), the optimal warpings are defined as: $\bm{\gamma}^{*}=\underset{\{\gamma_{i}\}}{\operatorname*{argmin}}Q_{\text{reg}}(\bm{\gamma})\triangleq\underset{\{\gamma_{i}\}}{\operatorname*{argmin}}{{\sum}}_{j=1}^{C}\frac{\sum\limits_{i\in\{y_{i}=j\}}\lVert Q_{i}(\bm{\gamma}_{i})-\bar{Q}^{(j)}\rVert}{N^{(j)}},\gamma_{i}\in\Gamma$, where “reg” means “registration”. For any other estimated warping $\hat{\bm{\gamma}}$, the registration error can be quantified as $\Delta Q_{\text{reg}}(\bm{\gamma}^{*},\hat{\bm{\gamma}})\triangleq|Q_{\text{reg}}(\bm{\gamma}^{*})-Q_{\text{reg}}(\hat{\bm{\gamma}})|.$
Theorem 3.1 (Low Registration Error).
Assume that: (i) The proposed deep neural network has sufficient capacity, (ii) Each warping function $\gamma(t)$ belongs to the admissible set $\Gamma$ and is continuously differentiable, and (iii) The SRVF of $q(t)$ of each curve $x(t)$ is continuous and bounded.
Then, for any $\epsilon>0$, there exists an estimated $\hat{\bm{\gamma}}$ produced by the model such that:
$\Delta Q_{\text{reg}}(\bm{\gamma}^{*},\hat{\bm{\gamma}})<\epsilon.$
Remark 3.2.
While Theorem 3.1 guarantees low registration error in theory, verifying this practically is challenging since the true warpings $\bm{\gamma}^{*}$ are not easily obtainable. This is due to the difficulty of achieving global optimization for $\min(Q_{\text{reg}}(\bm{\gamma}))$. However, the validity of the result can be assessed through simulation studies, where true warpings are provided.
We further show that the proposed model achieves a small generalization error. Let $\hat{f}_{\Theta}:\bm{x}\to y$ represent the mapping of the proposed DeepFRC model with a fixed architecture. Define the population risk and empirical risk as $R(\Theta)=\mathds{E}[l(\hat{f}_{\Theta}(\bm{x}),y)]$ and $R_{n}(\Theta)=\frac{1}{N}\sum_{i=1}^{N}l(\hat{f}_{\Theta}(\bm{x}_{i}),y_{i})$, respectively, where $l$ denotes the individual loss function. The generalization error is given by
$\Delta R_{\text{gen}}(\hat{\Theta})=|\mathds{E}_{S,A}[R(\hat{\Theta})-R_{n}(\hat{\Theta})]|,$
where $\hat{\Theta}$ is the parameter set estimated via a random algorithm $A$ based on a random sample $S$. During training, we assume the AdamW optimizer uses an inverse time decay or a custom decay function to ensure the learning rate $\alpha_{t}$ is monotonically non-increasing with $\alpha_{t}\leq c_{0}/t$, for some constant $c_{0}>0$, and that the algorithm runs for $T_{0}$ steps. The following result builds upon Theorem 3.8 from Hardt et al. (2016) and Theorem 2 from Yao et al. (2021).
Theorem 3.3 (Low Generalization Error).
Assume that: (i) There exists $\epsilon_{0}>0$, such that $|\bar{Q}^{(u)}-\bar{Q}^{(v)}|\geq\epsilon_{0}$ for $1\leq u<v\leq C$ and $\psi_{ij}\geq\epsilon_{0}$, and (ii) The weight $\Theta$ is restricted to a compact region.
Then, for the weights $\hat{\Theta}$ estimated by the proposed model, there exists a constant $c_{0}>0$, such that
$\Delta R_{\text{gen}}(\hat{\Theta})\lesssim\frac{T_{0}^{1-1/c_{0}}}{N}.$
Remark 3.4.
Assumption (i) in Theorem 3.3 is reasonable and easily satisfied. First, each class should exhibit a unique alignment pattern, meaning the corresponding $\bar{Q}$ values are distinct across classes. Second, by modifying the softmax function in the last layer of DeepFRC as: $\sigma(x_{j})=\frac{\text{exp}(x_{j})+\beta}{\sum_{i=1}^{C}(\text{exp}(x_{i})+\beta)}$, it is possible to select a constant $\beta$ such that each probability $\psi_{ij}$ is at least $\epsilon_{0}$.
4 Experiments
We present the results of functional data registration and classification using DeepFRC, evaluated on both simulated and real-world datasets. Simulated data provides insights into registration, reconstruction, and classification progression during training. Real-world datasets allow for performance comparisons with state-of-the-art methods. To test model generalization, training sets are deliberately smaller than test sets.
Evaluation Metrics.
We evaluate registration performance using the metric $\Delta Q_{\text{reg}}(\bm{\gamma}^{*},\hat{\bm{\gamma}})$ within the EFDA framework. This metric is applicable only to simulated datasets, as they provide true warping functions $\bm{\gamma}^{*}$ or underlying true curves. Another metric for alignment assessment is the adjusted total variance (ATV) defined as: $\text{ATV}=\frac{1}{\binom{C}{2}}\sum_{1\leq u<v\leq C}\frac{\text{TV}_{uv}}{d(\text{mean}(\tilde{x}_{u}),\text{mean}(\tilde{x}_{v}))},$ where $\text{TV}_{uv}$ (defined in (Chen and Srivastava, 2021)) is the total variation between classes $u$ and $v$, and $\text{mean}(\tilde{x})$ denotes the mean function of the aligned curve $\tilde{x}$. ATV accounts for the distance between class mean functions, making it more suitable for alignment evaluation of similar classes. Lower values of $\Delta Q_{\text{reg}}$ and ATV indicate better alignment. To assess DeepFRC’s ability to recover the underlying smooth process of aligned functional data, we compute the Pearson’s correlation coefficient $\rho(\bm{c}^{*},\hat{\bm{c}})$ between the true and estimated Fourier basis coefficients. This metric is used only in simulations, as $\bm{c}^{*}$ can only be computed when the true curves are known. Higher values of $\rho$ indicate better reconstruction. For classification performance, we evaluate accuracy (ACC) and $F_{1}$-score with the latter provides a more balanced measure of performance. We use the macro-averaged $F_{1}$-score (Sokolova and Lapalme, 2009) for multi-class classification. Higher values of ACC and $F_{1}$-score indicate better classification performance.
4.1 Simulation
Synthetic Data Generation.
We generate a balanced two-class functional dataset $\{x_{i}(t),y_{i}\}_{i=1}^{N}$, where each function $x_{i}(t)$ combines amplitude and phase variation. Amplitude is modeled by sums of Gaussian bumps, while phase variation is introduced via nonlinear warping functions $\gamma_{i}(t)$. Class-specific parameters and full generation details are provided in Appendix E.1. We simulate $N=6000$ functions, each evaluated at 1000 time points.
Results.
Figure 2 illustrates the effect of iterative optimization in DeepFRC on the simulated dataset. Figure 2(a) shows the raw curves of two classes (yellow and blue) with registration issues, which are challenging to distinguish visually. As shown in Figures 2(a)-(d), alignment improves progressively with the number of epochs, evidenced by the curves becoming more centered for each class, making them easier to differentiate. This process visually shows the mutual benefit of alignment and classification. Figure 2(e) demonstrates that the Pearson correlation coefficient $\rho(\bm{c}^{*},\hat{\bm{c}})$ increases rapidly over epochs, reaching nearly 1 by epoch 50, indicating DeepFRC’s strong reconstruction capability for the true unwarped functional data. Figures 2(f)-(g) show the progression of alignment and classification performance, all improving steadily with increasing epochs. Finally, Figure 2(h) presents the total loss function progression, which includes the registration loss (Eq. (10)) and classification loss (Eq. (11)). All three loss components decrease sharply before epoch 50 and stabilize near zero thereafter. Visualization of alignment and numerical results on the simulated test set are provided in Figure A1 of Appendix E.1, demonstrating excellent generalization performance.
Figure 2: Contribution of iterative optimization in training DeepFRC on simulated two-class (yellow and blue) data. (a)-(d) depict the progression of registration at Epochs 0, 5, 15, and 30. (e)-(h) show the improvement in reconstruction, registration, classification, and loss over SGD epochs.
4.2 Real Data Application
Real-World Datasets.
We evaluate our method on four publicly available datasets from the UCR Time Series Classification Archive, selected for their phase variation characteristics and common use in functional data analysis (Ramsay and Silverman, 2002, 2005): (1) Wave (motion): The Wave dataset contains 8 gesture classes recorded by accelerometers along three axes. We use the X-axis recordings from classes 2 and 8, yielding 1120 curves with 315 time points each. (2) Yoga (image-derived): The Yoga dataset features 1D time-series representations of yoga poses, derived by measuring the distance from the silhouette to the image center. It contains 3300 samples with 426 time points each. (3) Symbol (handwriting): The Symbol dataset records pen trajectories as participants replicate symbols. We use two subsets: a binary-class subset with 343 samples, and a three-class subset with 510 samples. Each sequence has 398 time points. The Wave and Symbol datasets are class-balanced, while Yoga is imbalanced. Further details are provided in Appendix E.2.
Baseline Models.
We benchmark DeepFRC against several deep learning models for functional data analysis, including a joint registration-classification model (TTN (Lohit et al., 2019)), a registration-only model (SrvfRegNet (Chen and Srivastava, 2021)), and four classification models: $\text{FCNN}_{\text{raw}}$, $\text{FCNN}_{\text{fourier}}$, FuncNN (Thind et al., 2020), and ADAFNN (Yao et al., 2021). To ensure a fair comparison, we evaluate sequential registration-classification pipelines by combining SrvfRegNet with each classification model (excluding TTN, which is already joint). A detailed description of all baselines appears in Appendix E.3.
Performance Comparison.
All models are trained to optimal performance; implementation details are given in Appendix E.4. Table 1 reports both alignment and classification metrics across four real-world datasets. Among all models, only DeepFRC and TTN jointly perform registration and classification. DeepFRC outperforms TTN by using elastic FDA for enhanced functional smoothness, delivering superior registration across all datasets and significant classification improvements on Symbol data. Unlike pipelines combining SrvfRegNet and classification models - which disregard label information during registration - DeepFRC achieves superior alignment, leading to more interpretable transformations and improved downstream classification accuracy. Figure A2 (Appendix E.4) plots the training loss versus epochs for DeepFRC across all datasets, confirming convergence.
Table 1: Quantitative comparison of registration and classification performance with state-of-the-art approaches across four real datasets. Bold indicates best results.
Wave Yoga Symbol (2 classes) Symbol (3 classes)
Model ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score
DeepFRC 5.6 96.4% 0.965 16.2 89.8% 0.909 4.8 96.0% 0.959 3.2 96.3% 0.963
TTN 6.3 94.7% 0.948 57.7 89.4% 0.904 8.6 92.0% 0.918 4.5 93.3% 0.933
SrvfRegNet+$\text{FCNN}_{\text{raw}}$ 7.3 94.6% 0.947 136.0 81.0% 0.830 14.8 94.5% 0.942 6.5 94.7% 0.947
SrvfRegNet+$\text{FCNN}_{\text{fourier}}$ 7.3 94.9% 0.950 136.0 84.0% 0.852 14.8 95.0% 0.949 6.5 96.0% 0.959
SrvfRegNet+FuncNN 7.3 95.7% 0.957 136.0 89.0% 0.908 14.8 93.5% 0.933 6.5 94.7% 0.947
SrvfRegNet+ADAFNN 7.3 94.6% 0.949 136.0 73.4% 0.753 14.8 89.0% 0.894 6.5 94.3% 0.943
Alignment Visualization.
Figure 3 compares alignment quality on the Symbol (3 classes) dataset. DeepFRC produces smooth, class-separated functional alignments, while TTN distorts class-specific trajectories (e.g., purple), and SrvfRegNet loses inter-class separation (e.g., blue vs. yellow). Both baselines significantly degrade interpretability (see also Figure A3, Appendix E.4). DeepFRC’s superior alignment improves inference reliability and preserves generalization performance - critical for real-world applications.
Figure 3: Visual comparison of alignment on the Symbol (3 classes).
Ablation Studies.
Table 2 presents ablation study results that separately evaluate the registration and classification components of our method. Removing the classification module (DeepFRC w/o Class.) degrades registration accuracy, while removing registration (DeepFRC w/o Reg.) harms classification - especially on Yoga and Symbol (2 classes). These results highlight the synergy between registration and classification within DeepFRC’s unified architecture.
Table 2: Ablation study: contributions of registration and classification components.
Wave Yoga Symbol (2 classes) Symbol (3 classes)
Model ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score
DeepFRC 5.6 96.4% 0.965 16.2 89.8% 0.909 4.8 96.0% 0.959 3.2 96.3% 0.963
DeepFRC w/o Class. 7.3 136.0 14.8 6.5
DeepFRC w/o Reg. 94.4% 0.946 83.1% 0.846 91.0% 0.905 94.7% 0.947
Sensitivity Analysis.
We assess DeepFRC’s robustness to the choice of basis family and size $K$ across four real datasets in Figure 4. The left panel shows performance is stable across basis types (Fourier, B-spline, polynomial). We attribute this robustness to two factors: (a) joint training adapts alignment to the inductive biases of each basis, and (b) class-aware registration loss guides the model toward discriminative structures, reducing basis-specific limitations. The right panel shows that increasing $K$ improves performance up to around $K=100$, after which it plateaus. We believe this invariance arises from (a) upstream registration reducing functional variability, and (b) downstream layers adaptively re-weighting or pruning redundant basis functions.
Figure 4: Left: Sensitivity to basis family. “F”: Fourier; “B”: B-spline; “P”: Polynomial. Right: Sensitivity to number of basis functions $K$.
5 Discussion
Robustness to Missing Data.
We evaluated DeepFRC’s tolerance to missing observations by randomly removing 5-10% of data points from half of each real-world dataset (Section 4.2), imputing gaps via Fourier splines (Wahba, 1990). Under identical hyperparameters, DeepFRC achieved comparable registration and classification performance to complete data (Figure A4, Appendix E.4 ), except on the noisy Yoga dataset, where alignment degraded slightly at 10% missingness. Results confirm robustness to minor missing data with proper smoothing.
Scalability to Large and Small Data.
DeepFRC scales efficiently to large datasets with $\mathcal{O}(Nn)$ complexity (Section 2.5). On the augmented Symbol dataset (100 $\times$), it achieves near-raw performance (ATV=4.8, $F_{1}$-score=0.944 vs. 4.8, 0.959). Simulations under sparse data conditions (samples, time points, or both) confirm robust classification, though registration requires sequence length $n\geq 100$ for reliability (Table A3, Appendix E.4).
6 Conclusion
This work introduces DeepFRC, a robust framework for joint functional registration and classification, supported by theoretical guarantees and extensive empirical validation. DeepFRC outperforms state-of-the-art methods and remains stable across real-world scenarios, including missing data and both small- and large-scale datasets. Assuming a shared hidden warping pattern across dimensions, DeepFRC naturally extends to $d$-dimensional functional data ($d\geq 2$) by replacing the input $x_{i}(\bm{t}_{i})\in\mathbb{R}^{n}$ with $x_{i}(\bm{t}_{i})\in\mathbb{R}^{n\times d}$, without introducing additional assumptions.
Broader Impacts and Limitations.
This is a methodological and theoretical contribution without direct societal implications. While DeepFRC exhibits empirical robustness to the choice and dimensionality of basis functions, a theoretical formalization of this invariance remains an open direction. Additionally, its performance may decline on highly volatile functional data, where the current objective is sensitive to rapid directional fluctuations. This suggests the need for more adaptive mechanisms or robust loss formulations to better accommodate extreme signal variability. Future work could incorporate signal complexity awareness into the model architecture or objective design to further improve generalization.
Appendix A Objective Function Gradient
By the chain rule, the gradient of the objective function with respect to any $\theta\in\Theta$ is derived as follows:
$\frac{\partial\mathcal{L}(\Theta)}{\partial\theta}=\begin{cases}2\sum\limits_{j=1}^{C}\frac{\sum\limits_{i\in\{y_{i}=j\}}\big{(}Q_{i}(\bm{\gamma}_{i})\big{)}^{\intercal}\cdot\frac{\partial Q_{i}(\bm{\gamma}_{i})}{\partial\gamma_{i}}\cdot\frac{\partial\gamma_{i}}{\partial\theta}}{N^{(j)}}-\frac{\beta}{N}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{C}\frac{y_{ij}}{\psi_{ij}}\cdot\frac{\partial\psi_{ij}}{\partial\theta},&\theta\in\Theta_{1},\\ -\frac{\beta}{N}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{C}\frac{y_{ij}}{\psi_{ij}}\cdot\frac{\partial\psi_{ij}}{\partial\theta},&\theta\in\Theta_{2}.\end{cases}$ (A.1)
Here, the gradient of the warped SRVF with respect to the warping function $\gamma_{i}$ is given by:
$\frac{\partial Q_{i}(\bm{\gamma}_{i})}{\partial\gamma_{i}}[k]=\dot{q}_{i}(\gamma_{i}(t_{ik}))\sqrt{\dot{\gamma}(t_{ik})}+q_{i}(\gamma_{i}(t_{ik}))\frac{\ddot{\gamma}(t_{ik})}{2\sqrt{\dot{\gamma}(t_{ik})}},\quad k=1,\dots,n.$ (A.2)
When computing the gradient of $\mathcal{L}_{1}(\Theta_{1})$ with respect to the warped SRVF ($Q_{i}(\bm{\gamma}_{i})$), the mean SRVF ($\bar{Q}^{(j)}$) is treated as constant, updated using the arithmetic mean of the warped SRVF from the previous iteration [Chen and Srivastava, 2021]. For $\theta\in\Theta_{1}$, the gradient $\frac{\partial\gamma_{i}}{\partial\theta}$ can be expressed as the product of gradients from integration, fully connected layers, and the 1D-CNN module. Similarly, $\frac{\partial\psi_{ij}}{\partial\theta}$ is derived from the gradient of $l_{2}$ fully connected layers, interpolation, integration, and the 1D-CNN module [Yann et al., 2015]. For $\theta\in\Theta_{2}$, $\frac{\partial\psi_{ij}}{\partial\theta}$ involves the gradient from fully connected layers in the prediction module [Yann et al., 2015].
Appendix B Model Selection of DeepFRC
The model’s hyperparameters include the size of basis functions $K$, loss-related parameters $\{\alpha,\beta\}$, and training-related parameters (such as epoch size $E$, batch size, and learning rate, denoted by vector $\bm{\eta}$). For the basis representation module, we set $K=100$ with a Fourier basis, and the sensitivity analysis will be studied in real data analysis. To select the other hyperparameters, we follow the procedure proposed by Wang et al. [2023]: the dataset $\{(x(\bm{t}_{i}),y_{i})\}_{i=1}^{N}$ is firstly split into two subsets with a 4:1 ratio. For each combination of hyperparameters, the model is trained by minimizing $\mathcal{L}_{\text{train}}$ (Eq. (12)) on the larger subset, and testing error $\mathcal{L}_{\text{test}}$ is computed on the smaller subset. The combination minimizing $\mathcal{L}_{\text{test}}$ is selected.
In the 1D-CNN module of our model, we fixed the setting of the network structure (set $l_{1}=4$, and set kernel sizes as 3-3-3 and the channel dimensions as 16-32-64 for the three hidden convolutional layers). To improve efficiency, robustness, and interpretability, we apply max-pooling [Nagi et al., 2011] and 1D batch normalization [Ioffe and Szegedy, 2015] after each layer to downsample inputs and stabilize training, and use global averaging [Krizhevsky et al., 2017] followed by the final fully connected layer to further reduce features and noise. In the classifier module, we set $l_{2}=3$ and nodes number $\bm{n}=(8,4)$. The above configuration for 1D-CNN module and the fully connected neural network classifier is widely used Chen and Srivastava [2021], Lohit et al. [2019], and works well across all the datasets in this paper.
Appendix C Proof of Theorem 3.1
For simplicity, we assume $\mathbf{t}_{i}=\left(\frac{0}{T},\cdots,\frac{T}{T}\right)$, where $[0,T]$ is the observed time range, in the following proof. Consider the function $f(x)=\sqrt{x}$, which is continuous on $[0,1]$. For any $\varepsilon>0$, there exists $\delta_{1}>0$ such that for all $|x_{1}-x_{2}|<\delta_{1}$, we have $\left|\sqrt{x_{1}}-\sqrt{x_{2}}\right|<\varepsilon.$ Similarly, since $q$ is continuous, for any $\varepsilon>0$, there exists $\delta_{2}>0$ such that for all $|x_{1}-x_{2}|<\delta_{2}$, we have $\left|q(x_{1})-q(x_{2})\right|<\varepsilon.$
Let $\mathscr{G}$ be the domain of the neural network’s output function, and $\Gamma$ be the domain of the alignment functions $\gamma$. Denote $\Phi:\mathscr{G}\to\Gamma$ as the mapping from the neural network output to the alignment functions. Since both input and output of the network are discrete, we consider values only at discrete points. We define $g_{i}^{(k)}=g_{i}\left(\frac{k}{T}\right)$ to maintain consistency between discrete vectors and continuous functions. The mapping is defined as:
$\tau_{i}\mapsto\tilde{\gamma}_{i}\mapsto\gamma_{i},$
$\tilde{\gamma}_{i}\left(\frac{k}{T}\right)=\frac{\sum_{s=0}^{k}\left(\tau_{i}^{(s)}\right)^{2}}{\sum_{s=0}^{T}\left(\tau_{i}^{(s)}\right)^{2}},\quad\gamma_{i}\left(\frac{k}{T}\right)=\frac{\sum_{s=0}^{k}\tilde{\gamma}_{i}^{(s)}}{\sum_{s=0}^{T}\tilde{\gamma}_{i}^{(s)}}.$
Let $\tau_{i}$ be such that $\Phi(\tau_{i})=\gamma_{i}$. Since $\Phi(t\tau_{i})=\Phi(\tau_{i})$ for all $t\neq 0$, we can assume, without loss of generality, that $\|g_{i}\|_{2}=1$.
By the approximation property of convolutional neural networks [Yarotsky, 2018] and classical universal approximation results [Cybenko, 1989, Funahashi, 1989, Hornik, 1991, Stinchcombe, 1999], for any $\delta_{4}>0$, there exists a neural network $NN_{\Theta}$ with parameters $\Theta$ such that
$\sup_{i}\|NN_{\Theta}(\mathbf{x}_{i})-\tau_{i}\|_{2}<\delta_{4}.$
Let $\hat{\tau}_{i}=NN_{\Theta}(\mathbf{x}_{i})$. We now aim to show that for sufficiently small $|\tau_{i}-\hat{\tau}_{i}|$, we also have $|\gamma_{i}-\hat{\gamma}_{i}|$ sufficiently small.
$\begin{split}\left|\tilde{\gamma}_{i}\left(\frac{k}{T}\right)-\hat{\tilde{\gamma}}_{i}\left(\frac{k}{T}\right)\right|&=\left|\frac{\sum_{s=0}^{k}\tau_{i}^{2}\left(\frac{s}{T}\right)}{\sum_{s=0}^{T}\tau_{i}^{2}\left(\frac{s}{T}\right)}-\frac{\sum_{s=0}^{k}\hat{\tau}_{i}^{2}\left(\frac{s}{T}\right)}{\sum_{s=0}^{T}\hat{\tau}_{i}^{2}\left(\frac{s}{T}\right)}\right|\\ &\leq\left|\frac{\sum_{s=0}^{k}\tau_{i}^{2}\left(\frac{s}{T}\right)}{\sum_{s=0}^{T}\tau_{i}^{2}\left(\frac{s}{T}\right)}-\frac{\sum_{s=0}^{k}\hat{\tau}_{i}^{2}\left(\frac{s}{T}\right)}{\sum_{s=0}^{T}\tau_{i}^{2}\left(\frac{s}{T}\right)}\right|+\left|\frac{\sum_{s=0}^{k}\hat{\tau}_{i}^{2}\left(\frac{s}{T}\right)}{\sum_{s=0}^{T}\tau_{i}^{2}\left(\frac{s}{T}\right)}-\frac{\sum_{s=0}^{k}\hat{\tau}_{i}^{2}\left(\frac{s}{T}\right)}{\sum_{s=0}^{T}\hat{\tau}_{i}^{2}\left(\frac{s}{T}\right)}\right|\\ &\leq\sum_{s=0}^{k}\left(\tau_{i}\left(\frac{s}{T}\right)+\hat{\tau}_{i}\left(\frac{s}{T}\right)\right)\delta_{4}+(1+\delta_{4})^{2}\frac{(2+\delta_{4})\delta_{4}}{(1-\delta_{4})^{2}}\\ &\leq\sum_{s=0}^{k}\left(\tau_{i}\left(\frac{s}{T}\right)+\hat{\tau}_{i}\left(\frac{s}{T}\right)\right)\delta_{4}+27\delta_{4}=M_{k}^{(1)}\delta_{4}.\end{split}$
Therefore, we have
$\|\tilde{\gamma}_{i}-\hat{\tilde{\gamma}}_{i}\|_{2}\leq\sqrt{\sum_{k=0}^{T}M_{k}^{1}}=M_{-1}^{(1)}\delta_{4}.$
By Hölder’s inequality $\|f\|_{2}\leq\|f\|_{1}\leq\sqrt{T+1}\|f\|_{2},$ we have
$\|\tilde{\gamma}_{i}-\hat{\tilde{\gamma}}_{i}\|_{1}\leq\sqrt{T+1}M_{-1}^{(1)}\delta_{4}.$
Note that $\|\tilde{\gamma}_{i}\|_{1}\geq\tilde{\gamma}_{i}\left(\frac{T}{T}\right)=1$. On the other hand, for $\gamma_{i}$, we have
$\begin{split}\left|\gamma_{i}\left(\frac{k}{T}\right)-\hat{\gamma}_{i}\left(\frac{k}{T}\right)\right|&=\left|\frac{\sum_{s=0}^{k}\tilde{\gamma}_{i}\left(\frac{s}{T}\right)}{\sum_{z=0}^{T}\tilde{\gamma}_{i}\left(\frac{z}{T}\right)}-\frac{\sum_{s=0}^{k}\hat{\tilde{\gamma}}_{i}\left(\frac{s}{T}\right)}{\sum_{z=0}^{T}\hat{\tilde{\gamma}}_{i}\left(\frac{z}{T}\right)}\right|\\ &=\left|\frac{\sum_{s=0}^{k}\tilde{\gamma}_{i}\left(\frac{s}{T}\right)}{\sum_{z=0}^{T}\tilde{\gamma}_{i}\left(\frac{z}{T}\right)}-\frac{\sum_{s=0}^{k}\tilde{\gamma}_{i}\left(\frac{s}{T}\right)}{\sum_{z=0}^{T}\hat{\tilde{\gamma}}_{i}\left(\frac{z}{T}\right)}+\frac{\sum_{s=0}^{k}\hat{\tilde{\gamma}}_{i}\left(\frac{s}{T}\right)}{\sum_{z=0}^{T}\hat{\tilde{\gamma}}_{i}\left(\frac{z}{T}\right)}-\frac{\sum_{s=0}^{k}\hat{\tilde{\gamma}}_{i}\left(\frac{s}{T}\right)}{\sum_{z=0}^{T}\tilde{\gamma}_{i}\left(\frac{z}{T}\right)}\right|\\ &\leq\left|\frac{\sum_{s=0}^{k}\tilde{\gamma}_{i}\left(\frac{s}{T}\right)}{\sum_{z=0}^{T}\tilde{\gamma}_{i}\left(\frac{z}{T}\right)}-\frac{\sum_{s=0}^{k}\hat{\tilde{\gamma}}_{i}\left(\frac{s}{T}\right)}{\sum_{z=0}^{T}\tilde{\gamma}_{i}\left(\frac{z}{T}\right)}\right|+\left|\frac{\sum_{s=0}^{k}\hat{\tilde{\gamma}}_{i}\left(\frac{s}{T}\right)}{\sum_{z=0}^{T}\tilde{\gamma}_{i}\left(\frac{z}{T}\right)}-\frac{\sum_{s=0}^{k}\hat{\tilde{\gamma}}_{i}\left(\frac{s}{T}\right)}{\sum_{z=0}^{T}\hat{\tilde{\gamma}}_{i}\left(\frac{z}{T}\right)}\right|\\ &\leq\left|\frac{\sum_{s=1}^{k}\left(\tilde{\gamma}_{i}\left(\frac{s}{T}\right)-\hat{\tilde{\gamma}}_{i}\left(\frac{s}{T}\right)\right)}{\sum_{z=0}^{T}\tilde{\gamma}_{i}\left(\frac{z}{T}\right)}\right|+\|\hat{\tilde{\gamma}}_{i}\|_{1}\left|\frac{\sum_{s=0}^{k}\left(\hat{\tilde{\gamma}}_{i}\left(\frac{s}{T}\right)-\tilde{\gamma}_{i}\left(\frac{s}{T}\right)\right)}{\left(\sum_{z=0}^{T}\tilde{\gamma}_{i}\left(\frac{z}{T}\right)\right)\left(\sum_{z=0}^{T}\hat{\tilde{\gamma}}_{i}\left(\frac{z}{T}\right)\right)}\right|\\ &\leq\frac{k\sqrt{T+1}M^{(1)}_{-1}\delta_{4}}{\|\tilde{\gamma}_{i}\|_{1}}+\left(\|\tilde{\gamma}_{i}\|_{1}+\sqrt{T+1}M^{(1)}_{-1}\delta_{4}\right)\frac{\delta_{4}}{\|h_{i}\|_{1}\left(\|\tilde{\gamma}_{i}\|_{1}-\sqrt{T+1}M^{(1)}_{-1}\delta_{4}\right)}\\ &\leq M_{k}^{(2)}\delta_{4}\quad\text{for some }M_{k}^{(2)}.\end{split}$
Such that we have
$\|\gamma_{i}-\hat{\gamma}_{i}\|_{2}\leq\sqrt{\sum\limits_{k=0}^{T}M_{k}^{(2)}}:=M_{-1}^{(2)}\delta_{4}.$
Recall that $Q_{i}(\gamma_{i})\left(\frac{k}{T}\right)=(q_{i}\circ\gamma_{i}\left(\frac{k}{T}\right))\cdot\sqrt{\dot{\gamma}_{i}\left(\frac{k}{T}\right)}$. Using the conditions that $q_{i}$ and $\gamma_{i}$ are bounded, we can bound the difference between $Q_{i}(\gamma_{i})$ and $Q_{i}(\hat{\gamma}_{i})$ by:
$\begin{split}&\left|Q_{i}(\gamma_{i})\left(\frac{k}{T}\right)-Q_{i}(\hat{\gamma}_{i})\left(\frac{k}{T}\right)\right|\\ &\leq\left|q_{i}\circ\gamma_{i}\left(\frac{k}{T}\right)\cdot\left(\sqrt{\dot{\gamma}_{i}\left(\frac{k}{T}\right)}-\sqrt{\dot{\hat{\gamma}}_{i}\left(\frac{k}{T}\right)}\right)\right|+\left|\sqrt{\dot{\hat{\gamma}}_{i}\left(\frac{k}{T}\right)}\cdot\left(q_{i}\circ\gamma_{i}\left(\frac{k}{T}\right)-q_{i}\circ\hat{\gamma}_{i}\left(\frac{k}{T}\right)\right)\right|\\ &\leq\sup_{t}|q_{i}(t)|\cdot\left|\sqrt{\dot{\gamma}_{i}\left(\frac{k}{T}\right)}-\sqrt{\dot{\hat{\gamma}}_{i}\left(\frac{k}{T}\right)}\right|+\sup_{t}\sqrt{\dot{\gamma}_{i}(t)}\cdot\left|q_{i}\left(\gamma_{i}\left(\frac{k}{T}\right)\right)-q_{i}\left(\hat{\gamma}_{i}\left(\frac{k}{T}\right)\right)\right|.\end{split}$
We use $\frac{\hat{\gamma}_{i}\left(\frac{k+1}{T}\right)-\hat{\gamma}_{i}\left(\frac{k-1}{T}\right)}{\frac{2}{T}}$ to approximate the derivative for $\hat{\gamma}_{i}$, so we have
$\begin{split}\left|\dot{\gamma}_{i}\left(\frac{k}{T}\right)-\dot{\hat{\gamma}}_{i}\left(\frac{k}{T}\right)\right|&\leq\left|\dot{\gamma}_{i}\left(\frac{k}{T}\right)-\frac{\gamma_{i}\left(\frac{k+1}{T}\right)-\gamma_{i}\left(\frac{k-1}{T}\right)}{\frac{2}{T}}\right|+\left|\frac{\gamma_{i}\left(\frac{k+1}{T}\right)-\gamma_{i}\left(\frac{k-1}{T}\right)}{\frac{2}{T}}-\frac{\hat{\gamma}_{i}\left(\frac{k+1}{T}\right)-\hat{\gamma}_{i}\left(\frac{k-1}{T}\right)}{\frac{2}{T}}\right|\\ &\leq\left|\dot{\gamma}_{i}\left(\frac{k}{T}\right)-\frac{\gamma_{i}\left(\frac{k+1}{T}\right)-\gamma_{i}\left(\frac{k-1}{T}\right)}{\frac{2}{T}}\right|+T\sup_{k}\left|\gamma_{i}\left(\frac{k}{T}\right)-\hat{\gamma}_{i}\left(\frac{k}{T}\right)\right|.\end{split}$
For $10<\delta_{3}<$ 1, there exists a $T>0$ such that:
$\left|\dot{\gamma}_{i}\left(\frac{k}{T}\right)-\frac{\gamma_{i}\left(\frac{k+1}{T}\right)-\gamma_{i}\left(\frac{k-1}{T}\right)}{\frac{2}{T}}\right|<\delta_{3},\quad\forall k.$
Here we can choose $\delta_{3}<\frac{\delta_{1}}{2},\ \delta_{2}=\min\{\delta_{2},\frac{\delta_{1}}{2T}\}$, and $\delta_{4}<\delta_{2}$. Then, we have
$\left|\sqrt{\dot{\gamma}_{i}\left(\frac{k}{T}\right)}-\sqrt{\dot{\hat{\gamma}}_{i}\left(\frac{k}{T}\right)}\right|<\varepsilon,\quad\left|q_{i}\circ\gamma_{i}\left(\frac{k}{T}\right)-q_{i}\circ\hat{\gamma}_{i}\left(\frac{k}{T}\right)\right|<\varepsilon,$
such that
$\left|Q_{i}(\gamma_{i})\left(\frac{k}{T}\right)-Q_{i}(\hat{\gamma}_{i})\left(\frac{k}{T}\right)\right|\leq M_{-1}^{(3)}\varepsilon,\quad\text{where }M_{-1}^{(3)}\text{ is a constant.}$.
This means at each fixed point $\frac{k}{T}$, the difference between $Q_{i}(\gamma_{i})$ and $Q_{i}(\hat{\gamma}_{i})$ is bounded by $M_{-1}^{(3)}\varepsilon$, and it is straightforward to have
$\|Q_{i}(\gamma_{i})-Q_{i}(\hat{\gamma}_{i})\|_{2}\leq\sqrt{T}M_{-1}^{3}\varepsilon.$
We finally get
$\begin{split}\Delta Q_{\text{reg}}(\gamma^{*},\hat{\gamma})&=\left|Q_{\text{reg}}(\gamma^{*})-Q_{\text{reg}}(\hat{\gamma})\right|\\ &=\left|\sum_{j=1}^{C}\frac{1}{N^{(j)}}\sum_{i\in\{y_{i}=j\}}\|Q_{i}(\gamma_{i}^{*})-\overline{Q}^{(j)}\|^{2}-\sum_{j=1}^{C}\frac{1}{N^{(j)}}\sum_{i\in\{y_{i}=j\}}\|Q_{i}(\hat{\gamma}_{i})-\overline{\hat{Q}}^{(j)}\|^{2}\right|\\ &\leq\sum_{j=1}^{C}\frac{1}{N^{(j)}}\sum_{i\in\{y_{i}=j\}}\left|\|Q_{i}(\gamma_{i}^{*})-\overline{Q}^{(j)}\|^{2}-\|Q_{i}(\hat{\gamma}_{i})-\overline{\hat{Q}}^{(j)}\|^{2}\right|.\end{split}$
We now simplify $\frac{1}{N^{(j)}}\sum_{i\in\{y_{i}=j\}}\left|\|Q_{i}(\gamma_{i}^{*})-\overline{Q}^{(j)}\|^{2}-\|Q_{i}(\hat{\gamma}_{i})-\overline{\hat{Q}}^{(j)}\|^{2}\right|$ as
$\frac{1}{N^{(j)}}\sum_{i\in\{y_{i}=j\}}\left|\|a_{i}\|_{2}^{2}+\|c\|_{2}^{2}+2\langle a_{i},b_{i}\rangle+2\langle b_{i},c\rangle+2\langle a_{i},c\rangle\right|,$
where $a_{i}=Q_{i}(\gamma_{i})-Q_{i}(\hat{\gamma}_{i}),b_{i}=Q_{i}(\hat{\gamma}_{i})-\overline{Q}^{(j)},c=\overline{Q}^{(j)}-\overline{\hat{Q}}^{(j)}$.
By the Cauchy-Schwarz inequality: $\langle x,y\rangle\leq\|x\|_{2}\|y\|_{2},$ we have $\|a_{i}\|_{2}\leq\sqrt{T}M_{-1}^{(3)}\varepsilon$, and $\|c\|_{2}=\left\|\frac{1}{N^{(j)}}\sum_{i\in\{y_{i}=j\}}a_{i}\right\|_{2}\leq\sqrt{T}M_{-1}^{(3)}\varepsilon$. We also have $\|b_{i}\|_{2}\leq M^{(4)}$ due to the boundness of $q_{i}$ and $\gamma_{i}$.
Thus, we have
$\left|\frac{1}{N^{(j)}}\sum_{i\in\{y_{i}=j\}}\|Q_{i}(\gamma_{i})-\overline{Q}^{(j)}\|_{2}^{2}-\frac{1}{N^{(j)}}\sum_{i\in\{y_{i}=j\}}\|Q_{i}(\hat{\gamma}_{i})-\overline{\hat{Q}}^{(j)}\|_{2}^{2}\right|\leq 4T\left(M_{-1}^{(3)}\right)^{2}\varepsilon^{2}+4\sqrt{T}M_{-1}^{3}M^{(4)}\varepsilon.$
Therefore, it is straightforward to have
$\Delta Q_{\text{reg}}(\gamma^{*},\hat{\gamma})\leq 4CT\left(M_{-1}^{(3)}\right)^{2}\varepsilon^{2}+4C\sqrt{T}M_{-1}^{(3)}M^{(4)}\varepsilon.$
Appendix D Proof of Theorem 3.3
Based on Theorem 2 from Yao et al. [2021], we only need to verify that the loss function and its gradient are both Lipschitz continuous. Since the mean SRVF ($\bar{Q}^{(j)}$’s) are treated as constants when computing their gradients, and we have the condition that there exists $\epsilon_{0}>0$, such that $|\bar{Q}^{(u)}-\bar{Q}^{(v)}|\geq\epsilon_{0}$ for $1\leq u<v\leq C$, we can omit the term $\sum_{1\leq u<v\leq C}\alpha\lVert\bar{Q}^{(u)}-\bar{Q}^{(v)}\rVert^{-1}$, and simplify the loss function of DeepFRC for any individual as
$l(\hat{f}_{\Theta}(\bm{x}_{i}),y_{i})=\sum_{j=1}^{C}\|Q_{i}-\overline{Q}^{(j)}\|\mathds{1}_{\{i\in\{y_{i}=j\}\}}-\sum_{j=1}^{C}y_{ij}\log\psi_{ij}.$
The loss function is obviously continuous for $\psi_{ij}\geq\epsilon_{0}$, as $\frac{1}{z}$ is continuous everywhere except at zero. Similar to Eqs. (A.1) and (A.2), the gradient of $l$ with respect to any $\theta\in\Theta$ can be easily derived. These gradients can be expressed as products of gradients from ReLU, max pooling, etc., which are Lipschitz continuous. Therefore, the theorem follows directly from Theorem 3.8 in Hardt et al. [2016].
Appendix E Experimental Details
E.1 Description of Simulated Datasets
We construct a balanced two-class functional dataset $\{x_{i}(t),y_{i}\}_{i=1}^{N}$, where each function $x_{i}(t)$ for $t\in[0,1]$ is generated by composing amplitude and phase components. The amplitude function is defined as the sum of two Gaussian bumps:
$z_{i}(t\,|\,a_{i},\mu,\sigma_{i})=a_{i}\cdot\exp\left(-\frac{1}{2}\frac{(t-\mu)^{2}}{\sigma_{i}^{2}}\right),$
where $a_{i}$ and $\sigma_{i}$ are independently drawn from uniform distributions, and $\mu$ is a fixed constant. To introduce phase variation, we define a warping function $\gamma_{i}(t)$ as
$\gamma_{i}(t)=\begin{cases}\frac{\exp(b_{i}t)-1}{\exp(b_{i})-1},&\text{if }b_{i}\neq 0,\\ t,&\text{if }b_{i}=0,\end{cases}\quad\text{with }b_{i}\sim U(-1.5,1.5).$
Such constructions are widely used for modeling functional data with phase variability [Kneip and Ramsay, 2008, Tucker et al., 2013]. We then generate each function as:
$x_{i}(t)=\begin{cases}\sum_{j=1}^{2}z_{i}(\gamma_{i}(t)\,|\,a_{i}^{(0j)},\mu^{(0j)},\sigma_{i}^{(0j)}),&\text{if }y_{i}=0,\\ \sum_{j=1}^{2}z_{i}(\gamma_{i}(t)\,|\,a_{i}^{(1j)},\mu^{(1j)},\sigma_{i}^{(1j)}),&\text{if }y_{i}=1,\end{cases}$
where the parameters $\{a_{i}^{(cj)},\mu^{(cj)},\sigma_{i}^{(cj)}\}_{j=1}^{2}$ for $c=0,1$ are summarized in in the following table. We split the balanced dataset into training, validation, and test sets with sizes 1600, 400, and 4000, respectively.
Table A1: Parameter setting for simulated data
Index (01) (02) (11) (12)
$a_{i}$ $13+U(-0.5,0.5)$ $12.5+U(-1,1)$ $12+U(-1,1)$ $13+U(-1.5,1.5)$
$\mu$ 0.250 0.715 0.225 0.695
$\sigma_{i}$ $0.06+U(-0.003,0.003)$ $0.075+U(-0.003,0.003)$ $0.06+U(-0.003,0.003)$ $0.1+U(-0.003,0.003)$
[…]
E.2 Description of Real Datasets
The Four time series datasets from the UCR Time Series Classification Archive (https://www.timeseriesclassification.com/index.php):
The Wave dataset [Liu and Yang, 2009] consists of eight simple gestures generated using accelerometers, collected through a particular procedure. For a participant, gestures are gathered when they hold a device and repeat certain gestures multiple times during a time period. The dataset includes X, Y, and Z dimensions with 8 classes. Here, we selected the X dimension and classes 2 and 8 for the experiment. There are 1120 samples with label partition 591/529, and the train/validation/test split is 320/80/720, with each sample having 315 time points.
The Yoga dataset consists of images of two actors (one male, one female) transitioning between yoga poses in front of a green screen. The task is to classify the images based on the actor’s gender. Each image was transformed into a one-dimensional series by measuring the distance from the outline of the actor to the center. The dataset contains 3300 samples with label partition 1770/1530, and the train/validation/test split is 800/200/2300. Each sample has 426 time points.
The Symbol dataset involves 13 participants, who were asked to replicate a randomly appearing symbol. There were 3 possible symbols, creating a total of 6 classes, and each participant made approximately 30 attempts. The dataset contains X-axis motion data recorded during the process of drawing the shapes. We conduct experiments on both the 2-class and 3-class cases using the Symbol dataset. In the 2-class case, there are 343 samples with label partition 182/167, and the train/validation/test split is 115/28/200. In the 3-class case, there are 510 samples with label partition 182/167/161, and the train/validation/test split is 168/42/300. Each sample has 398 time points.
E.3 Introduction of Baselines
Here we provide a brief introduction of the compared baseline methods:
  • TTN [Lohit et al., 2019]: A method for joint alignment and classification of discrete time series but does not account for curve smoothness.
  • SrvfRegNet [Chen and Srivastava, 2021]: An unsupervised method that aligns raw functional data via a 1D CNN.
  • $\text{FCNN}_{\text{raw}}$: A model that discretizes raw functional data into a vector, inputting it into a fully connected neural network.
  • $\text{FCNN}_{\text{fourier}}$: A model that transforms the functional data into Fourier basis coefficients and then feeds the resulting vector into a fully connected neural network.
  • FuncNN [Thind et al., 2020]: A model that inputs entire functional curves directly into a fully connected neural network.
  • ADAFNN [Yao et al., 2021]: A model that inputs entire functional curves, adapting the bases during learning.
E.4 Model Implementation Details on Real Data
For all real-world datasets, the functional input is Z-score standardized entry-wise based on the mean function and standard deviation. To handle missing values (such as NAs) in the functional input, we use the BasisSmoother function from the skfda.preprocessing.smoothing module in Python to impute missing values. All of these baseline models are trained until the best performance is achieved. All the experiments are conducted on NVIDIA GeForce RTX 3090 32G. The details of implementation for all models are provided as follows:
  • The DeepFRC network is implemented as a PyTorch neural network. The hyperparameter settings for DeepFRC are shown in Table A2.
    Table A2: Hyperparameter setting for DeepFRC on four real datasets
    Case $\alpha$ $\beta$ $lr_{\text{reg}}$ $lr_{\text{class}}$
    Wave 3000 50 1e-3 1e-3
    Yoga 600 15 1e-3 5e-4
    Symbol (2 classes) 300 15 1e-3 1e-3
    Symbol (3 classes) 10 10 1e-3 1e-3
    In the discussion of stability with respect to basis expansion for DeepFRC, we compute Fourier, B-spline, and Polynomial scores as follows: The Fourier scores are obtained using the fast Fourier transform function rfft from torch.nn.fft, extracting the first 100 scores. The B-spline scores are derived from the cubic B-spline basis with 98 knots uniformly distributed over the interval $[0,1]$, resulting in exactly 100 B-spline basis functions. The Polynomial scores are obtained using the Chebyshev polynomial orthogonal basis, taking the first 100 Chebyshev polynomial functions. The B-spline and Polynomial scores are both calculated using the BasisSmoother function from the skfda.preprocessing.smoothing module in Python.
  • The model SrvfRegNet processes functional data through a learnable pre-warping block with three 1D convolutional layers (16, 32, 64 channels, kernel size 3, ReLU/BatchNorm/pooling), followed by a fully connected linear layer to generate warping functions. It applies time warping by minimizing the SRVF loss under an MSE criterion. The input data is the full functional data requiring alignment. The output of this network is the time-warping function that can generate aligned functional data. We use the default architecture for the models SrvfRegNet.
  • Both $\text{FCNN}_{\text{raw}}$ and $\text{FCNN}_{\text{fourier}}$ have the same architecture: a 3-layer MLP, with the input layer consisting of $n$ observations or 100 Fourier scores (by default), followed by hidden layers with dimensions [16, 8], and finally outputting the classification results. Each layer is equipped with Layer Normalization, the ReLU activation function, and a dropout rate of 0.1. The input of $\text{FCNN}_{\text{raw}}$ is the raw data, while the input of $\text{FCNN}_{\text{fourier}}$ is derived from the first 100 Fourier coefficients of the functional data.
  • The models FuncNN and ADAFNN both take the entire curve as input. FuncNN uses a manually chosen basis expansion, while ADAFNN employs an adaptively learned basis layer for representing curve data. We use the default architecture for the models FuncNN and ADAFNN.
Figure A2: Training loss vs. epochs by DeepFRC across the four real-life datasets.
Figure A3: Visualization comparison of alignment performance by DeepFRC, TTN and SrvfRegNet across three two-class real datasets
Figure A4: Performance of DeepFRC with different missing rates for raw data across four real-world datasets.
Table A3: Robustness under data scarcity. $N$ and $n$ represent sample size and sequence length, respectively.
Sparse samples ($n=1000$) Sparse time points ($N=6000$) Sparse samples and time points $(N,n)$
Evaluation $N=200$ $N=100$ $N=50$ $n=200$ $n=100$ $n=50$ $(200,200)$ $(100,100)$ $(50,50)$
ATV 0.605 2.105 2.571 0.580 0.784 0.867 1.097 2.454 12.492
ACC 100.0% 99.3% 98.2% 100.0% 100.0% 100.0% 100.0% 95.0% 80.0%
$\textrm{F}_{1}$-score 1.000 0.992 0.980 1.000 1.000 1.000 1.000 0.947 0.750

### v3

@@
 Abstract
-Functional data - observations in the form of curves or trajectories - arise in diverse domains such as biomedical sensing, motion capture, and handwriting recognition.
-A core challenge in functional data analysis (FDA) is accounting for phase variability, where misaligned temporal patterns hinder accurate inference.
-We introduce DeepFRC, an end-to-end deep learning framework for joint functional registration and classification.
-Unlike conventional approaches that decouple alignment and prediction, DeepFRC integrates class-aware elastic warping and a learnable basis representation into a unified architecture.
-This design enables temporal alignment and dimensionality reduction to be jointly optimized with classification, improving both interpretability and accuracy.
-We establish the first theoretical connection between alignment quality and generalization error, and validate our model on synthetic and real-world benchmarks.
-DeepFRC consistently outperforms state-of-the-art methods, especially in scenarios with complex temporal misalignment.
-Code is available at: https://github.com/Drivergo-93589/DeepFRC.
+Functional data, representing curves or trajectories, are ubiquitous in fields like biomedicine and motion analysis.
+A fundamental challenge is phase variability—temporal misalignments that obscure underlying patterns and degrade model performance.
+Current methods often address registration (alignment) and classification as separate, sequential tasks.
+This paper introduces DeepFRC, an end-to-end deep learning framework that jointly learns diffeomorphic warping functions and a classifier within a unified architecture.
+DeepFRC combines a neural deformation operator for elastic alignment, a spectral representation using Fourier basis for smooth functional embedding, and a class-aware contrastive loss that promotes both intra-class coherence and inter-class separation.
+We provide the first theoretical guarantees for such a joint model, proving its ability to approximate optimal warpings and establishing a data-dependent generalization bound that formally links registration fidelity to classification performance.
+Extensive experiments on synthetic and real-world datasets demonstrate that DeepFRC consistently outperforms state-of-the-art methods in both alignment quality and classification accuracy, while ablation studies validate the synergy of its components.
+DeepFRC also shows notable robustness to noise, missing data, and varying dataset scales.
+Code is available at https://github.com/Drivergo-93589/DeepFRC.
 11footnotetext: † These authors contributed equally.22footnotetext: * Corresponding author: zengpch@shanghaitech.edu.cn
@@
 1.3 Contributions
-This work introduces a unified framework for functional registration and classification with the following key contributions: (1) Model Innovation: We propose (a) a learnable basis representation that preserves infinite-dimensional structure, (b) a class-aware registration loss tailored for multi-class alignment, and (c) a joint training strategy that enables mutual reinforcement between alignment and prediction - surpassing prior approaches (Lohit et al., 2019; Chen and Srivastava, 2021).
-(2) Theoretical Foundation: We establish the first theoretical connection between registration fidelity and classification generalization, offering new insights into how structured alignment impacts downstream learning.
-(3) Empirical Validation: Extensive experiments on both synthetic and real datasets demonstrate the model’s ability to jointly optimize registration and classification.
-DeepFRC consistently outperforms state-of-the-art baselines, with ablation studies validating the necessity of its integrated design.
-The paper follows the order of contributions outlined above.
+Our main contributions are: (1) Unified Architecture for Joint Learning.
+We introduce the first end-to-end deep learning model that integrates a neural deformation operator for diffeomorphic warping, a spectral representation for smooth functional encoding, and a class-aware contrastive loss, enabling mutual reinforcement between alignment and prediction.
+(2) Theoretical Foundations.
+We establish the first theoretical guarantees for such a model, proving approximation capabilities and providing a generalization bound linking registration fidelity to classification performance.
+(3) Extensive Empirical Validation.
+Extensive experiments on synthetic and real-world data demonstrate that DeepFRC consistently outperforms state-of-the-art methods in both tasks, with ablation studies confirming the synergy of its components and further analyses highlighting its robustness and computational efficiency.
 2 The Model
-Consider the data $\big{\{}(x_{i}(\bm{t}_{i}),y_{i})\big{\}}_{i=1}^{N}$, where $x_{i}(\bm{t}_{i})=\big{(}x_{i}(t_{i1}),\dots,x_{i}(t_{in})\big{)}^{\intercal}$ represents the observed values of functional data $x(t)$ at $n$ time points for the $i$-th sample, and $y_{i}\in\{1,\dots,C\}$ denotes its class label.
-The functional data $x(t)$ is assumed to be a mean-square continuous process defined over $t\in[0,1]$, and the observations are misaligned.
-We hypothesize the existence of an underlying nonlinear time warping map $\gamma:t\mapsto\tilde{t}$ and a prediction function $f:x(\tilde{t})\mapsto y$.
-Our goal is to estimate $\gamma$ and $f$ simultaneously by an end-to-end deep learning model.
-This allows us to algin new functional data $x^{*}(\bm{t}^{*})$ using $\gamma(\bm{t}^{*})$ and classify it with $y^{*}=f(x^{*}(\gamma(\bm{t}^{*})))$.
-In the following, we detail the development of our deep neural networks to approximate both mappings $\gamma$ and $f$.
-2.1 1D CNN for Time Warping
-Our approach firstly utilizes 1D convolutional neural networks (1D CNNs) to extract temporal features from functional data, capitalizing on their demonstrated effectiveness in capturing temporal patterns (Kiranyaz et al., 2016; Ince et al., 2016; Chen and Srivastava, 2021; Kiranyaz et al., 2021).
-These networks process observed functional data points to produce latent features, which are then used to construct time-warping functions, enabling better registration of functional data.
-Given an input sequence $x_{i}(\bm{t}_{i})\in\mathbb{R}^{n}$, we extract time-warping features $\tau(x_{i}(\bm{t}_{i}))\in\mathbb{R}^{n}$ using a multi-layer 1D CNN.
-Each layer applies a 1D convolution followed by a ReLU activation:
-$\bm{h}^{(l)}=\text{ReLU}(W^{(l)}\circledast\bm{h}^{(l-1)}+B^{(l)}),\quad l=1,\ldots,l_{1}-1,$ (1)
-where $\bm{h}^{(0)}=x_{i}(\bm{t}_{i})$. The final layer is a fully connected transformation:
-$\tau(x_{i}(\bm{t}_{i}))=\text{ReLU}(W^{(l_{1})}\bm{h}^{(l_{1}-1)}+\bm{b}^{(l_{1})}).$ (2)
-Each convolutional layer has learnable weights $W^{(l)}$ and biases $B^{(l)}$.
-Following Chen and Srivastava (2021), we set $l_{1}=4$ and use three hidden 1D‐CNN blocks (kernel size $3$; channels $16\!\rightarrow\!32\!\rightarrow\!64$).
-This encoder, parameterised by $\Theta_{1}$, outputs the warping field $\tau(x_{i}(\bm{t}_{i}))$ that captures temporal structure.
-To construct a boundary-preserving diffeomorphic warping function, the latent features $\tau(x_{i}(\bm{t}_{i}))$ are transformed as:
-$\tilde{\gamma}_{i}(t_{ij})=\frac{\sum_{\mu=0}^{j}\tau^{2}_{i\mu}}{\sum_{\nu=0}^{n}\tau^{2}_{i\nu}},\hskip 5.69054ptj=1,\dots,n,$ (3)
-where $t_{i0}=0$, $t_{in}=1,\tau_{i0}=0$, and $\tau_{i\mu}(\mu>0)$ is the $\mu$-th element of $\tau(x_{i}(\bm{t}_{i}))$.
-This enures that $\tilde{\gamma}_{i}(0)=0,\tilde{\gamma}_{i}(1)=1$ and $\frac{d(\tilde{\gamma}_{i}(t))}{dt}>0$, satisfying the requirements of a valid warping function (Chen and Srivastava, 2021).
-To reduce roughness and preserve the geometric structure, the warping function is further smoothed:
+We study supervised learning with functional data, where each sample consists of an observed trajectory and a categorical label.
+Formally, the dataset is $\{(x_{i}(\bm{t}_{i}),y_{i})\}_{i=1}^{N}$, where $x_{i}(\bm{t}_{i})=(x_{i}(t_{i1}),\ldots,x_{i}(t_{in}))^{\intercal}$ are observations of an underlying trajectory $x(t)$ sampled at irregular and potentially misaligned time points $\bm{t}_{i}=(t_{i1},\ldots,t_{in})$, and $y_{i}\in\{1,\dots,C\}$ is its class label.
+We view functional data as trajectories embedded in an infinite-dimensional space, where misalignment corresponds to latent time reparameterizations.
+Instead of treating alignment and prediction as two separate stages, we propose to jointly learn a diffeomorphic time-warping map $\gamma:t\mapsto\tilde{t}$ and a classifier $f:x(\tilde{t})\mapsto y$ in a single end-to-end architecture (Figure 1).
+This allows us to directly learn representations that are invariant to temporal misalignment while remaining discriminative for classification.
+Figure 1: Overview of DeepFRC.
+Multiple raw functional trajectories are first aligned via a neural deformation operator (1D CNN) that learns diffeomorphic time warping $\gamma(t)$, producing both warped curves and alignment functions.
+The aligned signals are then expanded in a Fourier basis to obtain spectral coefficients $c_{1},\dots,c_{K}$, which serve as inputs to a classifier network (MLP with Softmax output) for predicting class labels $y\in{1,\dots,C}$.
+Training is performed jointly by minimizing a contrastive geometric alignment loss and a classification loss.
+2.1 Neural Deformation Operator for Time Warping
+We introduce a neural deformation operator, parameterized by a 1D convolutional network (Kiranyaz et al., 2016; Ince et al., 2016; Kiranyaz et al., 2021), that learns velocity fields defining diffeomorphic warpings.
+This perspective differs from handcrafted registration methods: instead of explicitly constructing alignment rules, we parameterize $\gamma$ through a deep network optimized jointly with downstream prediction.
+Given an input sequence $x_{i}(\bm{t}_{i})\in\mathbb{R}^{n}$, a multi-layer 1D CNN extracts temporal features $\tau(x_{i}(\bm{t}_{i}))\in\mathbb{R}^{n}$:
+$\bm{h}^{(l)}=\mathrm{ReLU}(W^{(l)}\circledast\bm{h}^{(l-1)}+B^{(l)}),\quad l=1,\ldots,l_{1}-1,$ (1)
+with $\bm{h}^{(0)}=x_{i}(\bm{t}_{i})$.
+The final layer is fully connected: $\tau(x_{i}(\bm{t}_{i}))=\mathrm{ReLU}(W^{(l_{1})}\bm{h}^{(l_{1}-1)}+\bm{b}^{(l_{1})}).$ Following Chen and Srivastava (2021), we use $l_{1}=4$ convolutional blocks with kernel size $3$ and channels $16\!\rightarrow\!32\!\rightarrow\!64$.
+This encoder, parameterized by $\Theta_{1}$, produces latent features that encode temporal deformation fields.
+To construct a boundary-preserving diffeomorphism, we transform the latent features into a monotone cumulative sum: $\tilde{\gamma}_{i}(t_{ij})=\frac{\sum_{\mu=0}^{j}\tau^{2}_{i\mu}}{\sum_{\nu=0}^{n}\tau^{2}_{i\nu}},\quad j=1,\dots,n,$ where $t_{i0}=0,t_{in}=1$, and $\tau_{i0}=0$.
+This ensures $\tilde{\gamma}_{i}(0)=0,\tilde{\gamma}_{i}(1)=1$, and $\frac{d\tilde{\gamma}_{i}}{dt}>0$, satisfying diffeomorphic constraints (Chen and Srivastava, 2021).
+To further improve smoothness and geometric consistency, we apply an additional normalization:
 $\gamma_{i}(t_{ij})=\frac{\sum_{\mu=0}^{j}\tilde{\gamma}_{i\mu}}{\sum_{\nu=0}^{n}\tilde{\gamma}_{i\nu}}.$ (2)
-This smoothing step ensures a smooth and well-structured warping function (Chen and Srivastava, 2021).
-2.2 Basis Representation for Aligned Function
-By applying the time warping $\gamma_{i}(t_{ij})$ to the functional data $x_{i}(\bm{t}_{i})$, the aligned functional data $\tilde{x}_{i}(t)$ can be obtained using 1D linear interpolation on the points $\big{\{}\big{(}\gamma_{i}(t_{ij}),x_{i}(t_{ij})\big{)}\big{\}}_{j=1}^{n}$:
-$\begin{split}\tilde{x}_{i}(t)\triangleq\sum_{r=0}^{n-1}\Big{[}x(t_{ir})+\frac{x_{i}(t_{i(r+1)})-x_{i}(t_{ir})}{\gamma_{i}(t_{i(r+1)})-\gamma_{i}(t_{ir})}(t-\gamma_{i}(t_{ir}))\Big{]}\cdot\mathds{1}_{\{\gamma_{i}(t_{ir})\leq t\leq\gamma_{i}(t_{i(r+1)})\}},\end{split}$ (5)
-where $\mathds{1}_{\{\cdot\}}$ is the indicator function. We choose this linear interpolation because it is both stable and computationally efficient (Berrut and Trefethen, 2004).
-While traditional methods often discretize $\tilde{x}_{i}(t)$ into a high-dimensional vector $\tilde{\bm{x}}_{i}$ for input into neural networks, this approach suffers from the curse of dimensionality and fails to capture the smoothness of functional data, especially in noisy settings (Yao et al., 2021).
-To address these issues, we approximate $\tilde{x}_{i}(t)$ using basis expansion:
+This defines a neural operator that outputs valid, smooth, and data-adaptive warping functions.
+2.2 Spectral Representation of Aligned Functions
+Once trajectories are aligned by $\gamma_{i}$, we interpolate the aligned function $\tilde{x}_{i}(t)$ on points $\{(\gamma_{i}(t_{ij}),x_{i}(t_{ij}))\}_{j=1}^{n}$ using stable 1D linear interpolation:
+$\begin{split}\tilde{x}_{i}(t)\triangleq\sum_{r=0}^{n-1}\Big[x(t_{ir})+\frac{x_{i}(t_{i(r+1)})-x_{i}(t_{ir})}{\gamma_{i}(t_{i(r+1)})-\gamma_{i}(t_{ir})}(t-\gamma_{i}(t_{ir}))\Big]\cdot\mathds{1}_{\{\gamma_{i}(t_{ir})\leq t\leq\gamma_{i}(t_{i(r+1)})\}}.\end{split}$ (3)
+Rather than vectorizing $\tilde{x}_{i}(t)$ into a high-dimensional grid, which ignores smoothness and leads to inefficiency, we embed aligned functions into a compact spectral basis:
 $\tilde{x}_{i}(t)\approx\sum_{j=1}^{K}c_{ij}\phi_{j}(t),$ (4)
-where $\{\phi_{j}(t)\}_{j=1}^{K}$ are set to Fourier basis functions and $\{c_{ij}\}_{j=1}^{K}$ are coefficients to estimate.
-Using least squares, the coefficients $\bm{c}_{i}\triangleq(c_{i1},\dots,c_{iK})$ are estimated as:
-[…]
-[…]
-2.3 Fully Connected Neural Network Classifier
-To classify the $i$-th sample, we use a fully connected feedforward neural network that takes the input feature vector $\tilde{\bm{c}}_{i}\in\mathbb{R}^{K}$.
-The network has $l_{2}-1$ hidden layers with ReLU activations, followed by a softmax output layer:
-$\bm{h}^{(l)}=\text{ReLU}(W^{(l)}\bm{h}^{(l-1)}+\bm{b}^{(l)}),\quad l=l_{1}+1,\dots,l_{1}+l_{2}-1.$ (8)
-Here, $W^{(l)}$ and $\bm{b}^{(l)}$ are learnable weights and biases. The output $\psi(\tilde{\bm{c}}_{i})\in\mathbb{R}^{C}$ gives the class probabilities. The predicted label is:
-$\psi(\tilde{\bm{c}}_{i})=\text{Softmax}(W^{(l_{1}+l_{2})}\bm{h}^{(l_{1}+l_{2}-1)}+\bm{b}^{(l_{1}+l_{2})}).$ (9)
-The classifier is a three-layer MLP ($l_{2}=3$) with $8$ and $4$ hidden units, parameterised by $\Theta_{2}$.
-2.4 Objective Function Formulation
-We denote the overall parameters to estimate as $\Theta=\{\Theta_{1},\Theta_{2}\}$.
-Given training data $\{(x_{i}(\bm{t}_{i}),y_{i})\}_{i=1}^{N}$ (where $y_{i}\in\{1,\dots,C\}$), the objective function consists of two parts for estimating $\Theta$.
-Class-aware Registration Loss.
-We adopt the elastic Riemannian metric square-root velocity function (SRVF) (Srivastava and Klassen, 2016) for registration, which overcomes limitations of the $\mathds{L}^{2}$ metric such as lack of invariance to warping, asymmetry, and the pinching effect (Marron et al., 2015; Srivastava and Klassen, 2016).
-The SRVF of $x(t)$ is defined as $q(t)=\text{sign}(\dot{x}(t))\sqrt{|\dot{x}(t)|}$, and for the warped function $x(\gamma(t))$, it becomes $(q\star\gamma)(t)\triangleq q(\gamma(t))\sqrt{\dot{\gamma}(t)}$ (Joshi et al., 2007; Srivastava and Klassen, 2016).
-For the $i$-th sample, we denote the observed SRVF as $Q_{i}(\bm{\gamma}_{i})=\big{(}(q_{i}\star\gamma_{i})(t_{i1}),\dots,(q_{i}\star\gamma_{i})(t_{in})\big{)}^{\intercal}$.
-Inspired by the work of Chen and Srivastava (2021), we define the class-aware registration loss as:
-$\begin{split}\mathcal{L}_{1}(\Theta_{1})={{\sum}}\limits_{j=1}^{C}\frac{\sum\limits_{i\in\{y_{i}=j\}}\lVert Q_{i}(\bm{\gamma}_{i})-\bar{Q}^{(j)}\rVert}{N^{(j)}}+\alpha\sum\limits_{1\leq u<v\leq C}\lVert\bar{Q}^{(u)}-\bar{Q}^{(v)}\rVert^{-1},\end{split}$ (10)
-[…]
+where $\{\phi_{j}(t)\}_{j=1}^{K}$ are Fourier basis functions and $c_{ij}$ are coefficients estimated by least squares:
+$\tilde{\bm{c}}_{i}=\arg\!\min_{\{c_{ij}\}}\int_{0}^{1}\Big[\tilde{x}_{i}(t)-\sum_{j=1}^{K}c_{ij}\phi_{j}(t)\Big]^{2}\,dt=G^{-1}\bm{d}_{i},$ (5)
+with $G_{ij}=\int_{0}^{1}\phi_{i}(t)\phi_{j}(t)\,dt$ and $d_{ij}=\int_{0}^{1}\tilde{x}_{i}(t)\phi_{j}(t)\,dt$.
+This spectral representation, inspired by classical functional data analysis, acts as a form of neural Fourier features, yielding compact, smooth, and regularized embeddings of aligned functions.
+We employ 1D linear interpolation (Eq.
+(3)) for its numerical stability and computational efficiency, and a Fourier basis expansion (Eq.
+(4)) as it introduces no additional regularization hyperparameters.
+Crucially, both methods satisfy the Lipschitz continuity condition required by Theorem 3.3.
+2.3 Classifier Network
+The spectral coefficients $\tilde{\bm{c}}_{i}\in\mathbb{R}^{K}$ are fed into a fully connected classifier with $l_{2}-1$ hidden layers and ReLU activations:
+$\bm{h}^{(l)}=\mathrm{ReLU}(W^{(l)}\bm{h}^{(l-1)}+\bm{b}^{(l)}),\quad l=l_{1}+1,\ldots,l_{1}+l_{2}-1,$ (6)
+followed by a softmax layer: $\psi(\tilde{\bm{c}}_{i})=\mathrm{Softmax}(W^{(l_{1}+l_{2})}\bm{h}^{(l_{1}+l_{2}-1)}+\bm{b}^{(l_{1}+l_{2})}).$ We use a three-layer MLP ($l_{2}=3$) with hidden units of size $8$ and $4$, parameterized by $\Theta_{2}$.
+2.4 Objective: Contrastive Geometric Alignment and Classification
+We optimize the model with parameters $\Theta=\{\Theta_{1},\Theta_{2}\}$ by combining geometric alignment and discriminative classification.
+Class-aware Contrastive Alignment.
+To align functional trajectories while preserving class structure, we adopt the square-root velocity function (SRVF) representation $q(t)=\text{sign}(\dot{x}(t))\sqrt{|\dot{x}(t)|}$ (Srivastava and Klassen, 2016).
+For a warped trajectory $x(\gamma(t))$, the SRVF is $(q\star\gamma)(t)=q(\gamma(t))\sqrt{\dot{\gamma}(t)}$.
+For sample $i$, the observed SRVF vector is $Q_{i}(\bm{\gamma}_{i})=((q_{i}\star\gamma_{i})(t_{i1}),\ldots,(q_{i}\star\gamma_{i})(t_{in}))^{\intercal}$.
+We define a contrastive alignment loss:
+$\mathcal{L}_{1}(\Theta_{1})=\sum_{j=1}^{C}\frac{\sum_{i:y_{i}=j}\|Q_{i}(\bm{\gamma}_{i})-\bar{Q}^{(j)}\|}{N^{(j)}}+\alpha\sum_{1\leq u<v\leq C}\|\bar{Q}^{(u)}-\bar{Q}^{(v)}\|^{-1},$ (7)
+where $\bar{Q}^{(j)}$ is the class-wise SRVF mean.
+The first term encourages intra-class alignment; the second is a contrastive separation term that increases inter-class margins.
+Together, this yields a new contrastive–geometric alignment objective.
 Classification Loss.
-The cross-entropy loss is given by
-$\mathcal{L}_{2}(\Theta)=-\frac{1}{N}\sum\limits_{i=1}^{N}\sum_{j=1}^{C}y_{ij}\text{log}\psi_{ij},$ (11)
-where the label $y_{i}$ is represented as a one-hot encoded vector $(y_{i1},\dots,y_{iC})^{\intercal}$ where all elements are set to zero except for the element that corresponds to the active class, which is set to one.
-The combined objective is:
+The classifier is trained using standard cross-entropy:
+$\mathcal{L}_{2}(\Theta)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{C}y_{ij}\log\psi_{ij}.$ (8)
+Joint Objective.
+The full objective integrates both alignment and prediction:
 $\mathcal{L}(\Theta)=\mathcal{L}_{1}(\Theta_{1})+\beta\mathcal{L}_{2}(\Theta),$ (9)
-where $\beta$ balances registration and classification.
-Minimizing $\mathcal{L}_{1}$ aligns functions across classes while aiding classification, especially for similar classes, while minimizing $\mathcal{L}_{2}$ predicts class labels and improves alignment by enhancing class separation.
-Integrating Subsections 2.1-2.4, we propose an end-to-end deep learning model, termed DeepFRC, for simultaneous functional registration and classification (Figure 1).
-Figure 1: Overview of DeepFRC.
-The light blue region highlights the 1D convolutional module.
-The three blue arrows (left to right) denote: flattening of the $(l_{1}{-}1)$th layer matrix, the time warping (TW) operation, and the basis projection of the aligned function.
-“TW” and “DR” denote time warping and dimensionality reduction, respectively.
-We set $l_{1}=4,l_{2}=3$.
+where $\beta$ balances geometric alignment and classification accuracy.
+Summary.
+Integrating Sections 2.12.4, we obtain DeepFRC (Deep Functional Registration and Classification), an end-to-end framework coupling (i) a neural deformation operator for alignment, (ii) a spectral embedding for smooth functional representation, and (iii) a classifier guided by a contrastive–geometric loss (Figure 1).
+To the best of our knowledge, DeepFRC is the first model to unify alignment and prediction of functional data through diffeomorphic neural operators and contrastive geometry, offering a new paradigm for learning invariant representations of misaligned functional trajectories.
+A natural extension of our framework allows it to handle $d$-dimensional functional inputs ($d\geq 2$) by simply considering $x_{i}(\bm{t}_{i})\in\mathbb{R}^{n\times d}$, under the assumption of a shared warping process across dimensions, without requiring additional structural assumptions.
 2.5 Optimization, Model Selection and Computational Complexity
@@
 Detailed proofs are provided in Appendices C and D.
-The phase variability is modeled using monotone and smooth transformations within the set $\Gamma=\{\gamma:[0,1]\to[0,1]\mid\gamma(0)=0,\gamma(1)=1,\dot{\gamma}>0\}$.
-Under the framework of elastic functional data analysis (EFDA) (Srivastava et al., 2011), the optimal warpings are defined as: $\bm{\gamma}^{*}=\underset{\{\gamma_{i}\}}{\operatorname*{argmin}}Q_{\text{reg}}(\bm{\gamma})\triangleq\underset{\{\gamma_{i}\}}{\operatorname*{argmin}}{{\sum}}_{j=1}^{C}\frac{\sum\limits_{i\in\{y_{i}=j\}}\lVert Q_{i}(\bm{\gamma}_{i})-\bar{Q}^{(j)}\rVert}{N^{(j)}},\gamma_{i}\in\Gamma$, where “reg” means “registration”.
-For any other estimated warping $\hat{\bm{\gamma}}$, the registration error can be quantified as $\Delta Q_{\text{reg}}(\bm{\gamma}^{*},\hat{\bm{\gamma}})\triangleq|Q_{\text{reg}}(\bm{\gamma}^{*})-Q_{\text{reg}}(\hat{\bm{\gamma}})|.$
+[…]
 Theorem 3.1 (Low Registration Error).
@@
 Remark 3.2.
-While Theorem 3.1 guarantees low registration error in theory, verifying this practically is challenging since the true warpings $\bm{\gamma}^{*}$ are not easily obtainable.
-This is due to the difficulty of achieving global optimization for $\min(Q_{\text{reg}}(\bm{\gamma}))$.
-However, the validity of the result can be assessed through simulation studies, where true warpings are provided.
+Theorem 3.1 provides a theoretical foundation for our approach, establishing that neural networks can approximate smooth warping functions in SRVF space, thereby justifying the use of a learnable diffeomorphic registration module.
+While direct empirical verification of the approximation error is challenging due to the inability to obtain globally optimal warpings $\bm{\gamma}^{*}$ in real data, we validate the practical observability of this result through controlled simulation studies where the ground-truth warpings are known.
 We further show that the proposed model achieves a small generalization error.
@@
 The generalization error is given by
-$\Delta R_{\text{gen}}(\hat{\Theta})=|\mathds{E}_{S,A}[R(\hat{\Theta})-R_{n}(\hat{\Theta})]|,$
+[…]
 where $\hat{\Theta}$ is the parameter set estimated via a random algorithm $A$ based on a random sample $S$.
@@
 Remark 3.4.
-Assumption (i) in Theorem 3.3 is reasonable and easily satisfied.
-First, each class should exhibit a unique alignment pattern, meaning the corresponding $\bar{Q}$ values are distinct across classes.
-Second, by modifying the softmax function in the last layer of DeepFRC as: $\sigma(x_{j})=\frac{\text{exp}(x_{j})+\beta}{\sum_{i=1}^{C}(\text{exp}(x_{i})+\beta)}$, it is possible to select a constant $\beta$ such that each probability $\psi_{ij}$ is at least $\epsilon_{0}$.
+The assumptions of Theorem 3.3 are empirically satisfied and guide model design.
+Class separation $|\bar{Q}^{(u)}-\bar{Q}^{(v)}|\geq\epsilon_{0}$ is supported by distinct SRVF means (Figure 3, Table 1).
+The probability floor $\psi_{ij}\geq\epsilon_{0}$ is enforced via additively smoothed softmax, keeping all probabilities above $10^{-4}$.
+Although non-constructive, the bound informs hyperparameter tuning: $\alpha$ is chosen to maximize separation, while $\beta$ balances classification and alignment.
 4 Experiments
@@
 Real-world datasets allow for performance comparisons with state-of-the-art methods.
-To test model generalization, training sets are deliberately smaller than test sets.
 Evaluation Metrics.
-We evaluate registration performance using the metric $\Delta Q_{\text{reg}}(\bm{\gamma}^{*},\hat{\bm{\gamma}})$ within the EFDA framework.
-This metric is applicable only to simulated datasets, as they provide true warping functions $\bm{\gamma}^{*}$ or underlying true curves.
-Another metric for alignment assessment is the adjusted total variance (ATV) defined as: $\text{ATV}=\frac{1}{\binom{C}{2}}\sum_{1\leq u<v\leq C}\frac{\text{TV}_{uv}}{d(\text{mean}(\tilde{x}_{u}),\text{mean}(\tilde{x}_{v}))},$ where $\text{TV}_{uv}$ (defined in (Chen and Srivastava, 2021)) is the total variation between classes $u$ and $v$, and $\text{mean}(\tilde{x})$ denotes the mean function of the aligned curve $\tilde{x}$.
-ATV accounts for the distance between class mean functions, making it more suitable for alignment evaluation of similar classes.
-Lower values of $\Delta Q_{\text{reg}}$ and ATV indicate better alignment.
-To assess DeepFRC’s ability to recover the underlying smooth process of aligned functional data, we compute the Pearson’s correlation coefficient $\rho(\bm{c}^{*},\hat{\bm{c}})$ between the true and estimated Fourier basis coefficients.
-This metric is used only in simulations, as $\bm{c}^{*}$ can only be computed when the true curves are known.
-Higher values of $\rho$ indicate better reconstruction.
-For classification performance, we evaluate accuracy (ACC) and $F_{1}$-score with the latter provides a more balanced measure of performance.
-We use the macro-averaged $F_{1}$-score (Sokolova and Lapalme, 2009) for multi-class classification.
-Higher values of ACC and $F_{1}$-score indicate better classification performance.
+We evaluate our method using metrics for registration, reconstruction, and classification.
+Registration quality is assessed by the alignment error ($\Delta Q_{\text{reg}}$) between estimated and true warping functions on simulated data, and by the Adjusted Total Variance (ATV) (Chen and Srivastava, 2021) on both simulated and real data, where lower values indicate better alignment.
+The fidelity of the recovered smooth process is measured via the correlation between true and estimated basis coefficients (simulated data only).
+Classification performance is reported using accuracy (ACC) and the macro-averaged $F_{1}$-score (Sokolova and Lapalme, 2009), with higher values being better.
+Detailed metric definitions are provided in Appendix E.1.
 4.1 Simulation
@@
 Class-specific parameters and full generation details are provided in Appendix E.2.
-We simulate $N=6000$ functions, each evaluated at 1000 time points.
 Results.
-Figure 2 illustrates the effect of iterative optimization in DeepFRC on the simulated dataset.
-Figure 2(a) shows the raw curves of two classes (yellow and blue) with registration issues, which are challenging to distinguish visually.
-As shown in Figures 2(a)-(d), alignment improves progressively with the number of epochs, evidenced by the curves becoming more centered for each class, making them easier to differentiate.
-This process visually shows the mutual benefit of alignment and classification.
-Figure 2(e) demonstrates that the Pearson correlation coefficient $\rho(\bm{c}^{*},\hat{\bm{c}})$ increases rapidly over epochs, reaching nearly 1 by epoch 50, indicating DeepFRC’s strong reconstruction capability for the true unwarped functional data.
-Figures 2(f)-(g) show the progression of alignment and classification performance, all improving steadily with increasing epochs.
-Finally, Figure 2(h) presents the total loss function progression, which includes the registration loss (Eq.
-(10)) and classification loss (Eq.
-(11)).
-All three loss components decrease sharply before epoch 50 and stabilize near zero thereafter.
-Visualization of alignment and numerical results on the simulated test set are provided in Figure A1 of Appendix E.1, demonstrating excellent generalization performance.
+[…]
 Figure 2: Contribution of iterative optimization in training DeepFRC on simulated two-class (yellow and blue) data.
@@
 Real-World Datasets.
-We evaluate our method on four publicly available datasets from the UCR Time Series Classification Archive, selected for their phase variation characteristics and common use in functional data analysis (Ramsay and Silverman, 2002, 2005): (1) Wave (motion): The Wave dataset contains 8 gesture classes recorded by accelerometers along three axes.
-We use the X-axis recordings from classes 2 and 8, yielding 1120 curves with 315 time points each.
-(2) Yoga (image-derived): The Yoga dataset features 1D time-series representations of yoga poses, derived by measuring the distance from the silhouette to the image center.
-It contains 3300 samples with 426 time points each.
-(3) Symbol (handwriting): The Symbol dataset records pen trajectories as participants replicate symbols.
-We use two subsets: a binary-class subset with 343 samples, and a three-class subset with 510 samples.
-Each sequence has 398 time points.
-The Wave and Symbol datasets are class-balanced, while Yoga is imbalanced.
+We evaluate our method on five publicly available datasets, selected for their prevalence in functional data analysis (Ramsay and Silverman, 2002, 2005) and relevance to phase variation: Wave, Yoga, Symbol, and MotionSense (Malekzadeh et al., 2019).
+Wave, Yoga, and Symbol are sourced from the UCR Time Series Classification Archive, while MotionSense is from Kaggle.
+For the Symbol dataset, we use both binary and three-class subsets.
+The Wave, Symbol, and MotionSense datasets are class-balanced, whereas Yoga is imbalanced.
+In terms of dimensionality, Wave, Yoga, and Symbol are one-dimensional, and MotionSense is three-dimensional.
 Further details are provided in Appendix E.3.
 Baseline Models.
-We benchmark DeepFRC against several deep learning models for functional data analysis, including a joint registration-classification model (TTN (Lohit et al., 2019)), a registration-only model (SrvfRegNet (Chen and Srivastava, 2021)), and four classification models: $\text{FCNN}_{\text{raw}}$, $\text{FCNN}_{\text{fourier}}$, FuncNN (Thind et al., 2020), and ADAFNN (Yao et al., 2021).
+We benchmark DeepFRC against several deep learning models for functional data analysis, including a joint registration-classification model (TTN (Lohit et al., 2019)), a registration-only model (SrvfRegNet (Chen and Srivastava, 2021)), and five classification models: $\text{FCNN}_{\text{raw}}$, $\text{FCNN}_{\text{fourier}}$, FuncNN (Thind et al., 2020), ADAFNN (Yao et al., 2021), and TSLANet (Eldele et al., 2024).
 To ensure a fair comparison, we evaluate sequential registration-classification pipelines by combining SrvfRegNet with each classification model (excluding TTN, which is already joint).
-A detailed description of all baselines appears in Appendix E.3.
+A detailed description of all baseline methods and implementation details are provided in Appendices E.4 and E.5, respectively.
 Performance Comparison.
-All models are trained to optimal performance; implementation details are given in Appendix E.4.
-Table 1 reports both alignment and classification metrics across four real-world datasets.
+Table 1 reports both alignment and classification metrics across five real-world datasets.
 Among all models, only DeepFRC and TTN jointly perform registration and classification.
-DeepFRC outperforms TTN by using elastic FDA for enhanced functional smoothness, delivering superior registration across all datasets and significant classification improvements on Symbol data.
+DeepFRC outperforms TTN by using elastic FDA for enhanced functional smoothness, delivering superior registration across all datasets and significant classification improvements on Symbol and MotionSense data.
 Unlike pipelines combining SrvfRegNet and classification models - which disregard label information during registration - DeepFRC achieves superior alignment, leading to more interpretable transformations and improved downstream classification accuracy.
+We observe that DeepFRC achieves classification accuracy comparable to TSLANet, a recent state-of-the-art time-series model that outperforms many transformer-based alternatives, across all datasets.
 Figure A2 (Appendix E.5) plots the training loss versus epochs for DeepFRC across all datasets, confirming convergence.
-Table 1: Quantitative comparison of registration and classification performance with state-of-the-art approaches across four real datasets. Bold indicates best results.
-Wave Yoga Symbol (2 classes) Symbol (3 classes)
-Model ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score
-DeepFRC 5.6 96.4% 0.965 16.2 89.8% 0.909 4.8 96.0% 0.959 3.2 96.3% 0.963
-TTN 6.3 94.7% 0.948 57.7 89.4% 0.904 8.6 92.0% 0.918 4.5 93.3% 0.933
-SrvfRegNet+$\text{FCNN}_{\text{raw}}$ 7.3 94.6% 0.947 136.0 81.0% 0.830 14.8 94.5% 0.942 6.5 94.7% 0.947
-SrvfRegNet+$\text{FCNN}_{\text{fourier}}$ 7.3 94.9% 0.950 136.0 84.0% 0.852 14.8 95.0% 0.949 6.5 96.0% 0.959
-SrvfRegNet+FuncNN 7.3 95.7% 0.957 136.0 89.0% 0.908 14.8 93.5% 0.933 6.5 94.7% 0.947
-SrvfRegNet+ADAFNN 7.3 94.6% 0.949 136.0 73.4% 0.753 14.8 89.0% 0.894 6.5 94.3% 0.943
+Table 1: Quantitative comparison of registration and classification performance with state-of-the-art approaches across five real datasets. Bold indicates best results.
+Wave Yoga Symbol (2 classes) Symbol (3 classes) MotionSense
+Model ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score
+DeepFRC 5.6 96.4% 0.965 16.2 89.8% 0.909 4.8 96.0% 0.959 3.2 96.3% 0.963 25.0 95.0% 0.952
+TTN 6.3 94.7% 0.948 57.7 89.4% 0.904 8.6 92.0% 0.918 4.5 93.3% 0.933 35.1 85.0% 0.857
+[…]
+[…]
+SrvfRegNet+FuncNN 7.3 95.7% 0.957 136.0 89.0% 0.908 14.8 93.5% 0.933 6.5 94.7% 0.947 37.7 90.0% 0.889
+SrvfRegNet+ADAFNN 7.3 94.6% 0.949 136.0 73.4% 0.753 14.8 89.0% 0.894 6.5 94.3% 0.943 37.7 85.0% 0.857
+SrvfRegNet+TSLANet 7.3 96.4% 0.961 136.0 89.3% 0.884 14.8 95.5% 0.955 6.5 96.3% 0.960 37.7 95.0% 0.952
 Alignment Visualization.
@@
 yellow).
-Both baselines significantly degrade interpretability (see also Figure A3, Appendix E.4).
+Both baselines significantly degrade interpretability (see also Figures A3 and A4, Appendix E.5).
 DeepFRC’s superior alignment improves inference reliability and preserves generalization performance - critical for real-world applications.
@@
 Ablation Studies.
-Table 2 presents ablation study results that separately evaluate the registration and classification components of our method.
-Removing the classification module (DeepFRC w/o Class.) degrades registration accuracy, while removing registration (DeepFRC w/o Reg.) harms classification - especially on Yoga and Symbol (2 classes).
-These results highlight the synergy between registration and classification within DeepFRC’s unified architecture.
-Table 2: Ablation study: contributions of registration and classification components.
-Wave Yoga Symbol (2 classes) Symbol (3 classes)
-Model ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score
-DeepFRC 5.6 96.4% 0.965 16.2 89.8% 0.909 4.8 96.0% 0.959 3.2 96.3% 0.963
-DeepFRC w/o Class. 7.3 136.0 14.8 6.5
-DeepFRC w/o Reg. 94.4% 0.946 83.1% 0.846 91.0% 0.905 94.7% 0.947
+The ablation study in Table 2 evaluates the individual contributions of the core components of our method: the Neural Deformation Operator (N.D.O.) for registration, the Spectral Representation (S.R.), and the Classifier Network (C.N.).
+The results demonstrate a clear synergy between these modules.
+Removing the registration component (DeepFRC w/o N.D.O.) significantly degrades classification performance, particularly on the Yoga, Symbol (2-class), and MotionSense datasets.
+Conversely, removing the classifier (DeepFRC w/o C.N.) impairs registration accuracy.
+The removal of the spectral representation (DeepFRC w/o S.R.) adversely affects both tasks across all datasets.
+These findings underscore the importance of the spectral representation and the tight coupling between registration and classification within DeepFRC’s unified architecture.
+Table 2: Ablation study: contributions of three components in DeepFRC.
+Wave Yoga Symbol (2 classes) Symbol (3 classes) MotionSense
+Model ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score
+DeepFRC 5.6 96.4% 0.965 16.2 89.8% 0.909 4.8 96.0% 0.959 3.2 96.3% 0.963 25.0 95.0% 0.952
+DeepFRC w/o N.D.O. 94.4% 0.946 83.1% 0.846 91.0% 0.905 94.7% 0.947 90.0% 0.909
+DeepFRC w/o S.R. 5.8 95.3% 0.955 17.7 89.2% 0.903 5.3 94.5% 0.945 3.3 93.3% 0.933 28.8 90.0% 0.900
+DeepFRC w/o C.N. 7.3 136.0 14.8 6.5 37.7
 Sensitivity Analysis.
-We assess DeepFRC’s robustness to the choice of basis family and size $K$ across four real datasets in Figure 4.
+We assess DeepFRC’s robustness to the choice of basis family and size $K$ across five real datasets in Figure 4.
 The left panel shows performance is stable across basis types (Fourier, B-spline, polynomial).
@@
 5 Discussion
-Robustness to Missing Data.
-We evaluated DeepFRC’s tolerance to missing observations by randomly removing 5-10% of data points from half of each real-world dataset (Section 4.2), imputing gaps via Fourier splines (Wahba, 1990).
-Under identical hyperparameters, DeepFRC achieved comparable registration and classification performance to complete data (Figure A4, Appendix E.4 ), except on the noisy Yoga dataset, where alignment degraded slightly at 10% missingness.
-Results confirm robustness to minor missing data with proper smoothing.
+Computational Efficiency.
+DeepFRC offers significant computational advantages during inference, with an average runtime of 60s on our real-world datasets—dramatically faster than the traditional registration method DTW (1000s) and competitive with deep baselines like TTN (15s) and SrvfRegNet hybrids (90–100s).
+This efficiency is achieved by eschewing pairwise dynamic programming.
+Although training time is on par with other deep models, DeepFRC achieves superior accuracy, striking an excellent trade-off between performance and computational cost.
+Robustness to Noise and Reconstruction Accuracy.
+We evaluate robustness by injecting Gaussian noise ($\sigma=0,0.05,0.10,0.20$) into synthetic data and measuring reconstruction quality via the Pearson correlation between true and estimated basis coefficients.
+As shown in Table A3, DeepFRC maintains high correlation under low-to-moderate noise, outperforming TTN and exhibiting greater stability than SrvfRegNet.
+Even under high noise ($\sigma=0.20$), its performance degrades gracefully while surpassing competitors, indicating that the smooth basis representation and joint optimization confer strong resilience.
 Scalability to Large and Small Data.
@@
 Simulations under sparse data conditions (samples, time points, or both) confirm robust classification, though registration requires sequence length $n\geq 100$ for reliability (Table A4, Appendix E.5).
+Robustness to Missing Data.
+We evaluated DeepFRC’s tolerance to missing observations by randomly removing 5-10% of data points from half of each real-world dataset (Section 4.2), imputing gaps via Fourier splines (Wahba, 1990).
+Under identical hyperparameters, DeepFRC achieved comparable registration and classification performance to complete data (Figure A5, Appendix E.5 ).
 6 Conclusion
-This work introduces DeepFRC, a robust framework for joint functional registration and classification, supported by theoretical guarantees and extensive empirical validation.
-DeepFRC outperforms state-of-the-art methods and remains stable across real-world scenarios, including missing data and both small- and large-scale datasets.
-Assuming a shared hidden warping pattern across dimensions, DeepFRC naturally extends to $d$-dimensional functional data ($d\geq 2$) by replacing the input $x_{i}(\bm{t}_{i})\in\mathbb{R}^{n}$ with $x_{i}(\bm{t}_{i})\in\mathbb{R}^{n\times d}$, without introducing additional assumptions.
-Broader Impacts and Limitations.
-This is a methodological and theoretical contribution without direct societal implications.
-While DeepFRC exhibits empirical robustness to the choice and dimensionality of basis functions, a theoretical formalization of this invariance remains an open direction.
-Additionally, its performance may decline on highly volatile functional data, where the current objective is sensitive to rapid directional fluctuations.
-This suggests the need for more adaptive mechanisms or robust loss formulations to better accommodate extreme signal variability.
-Future work could incorporate signal complexity awareness into the model architecture or objective design to further improve generalization.
+We introduced DeepFRC, a unified framework for joint functional registration and classification with theoretical guarantees and comprehensive empirical validation.
+DeepFRC consistently outperforms state-of-the-art baselines and remains robust across real-world and simulated settings with noise, missing values, and scale variation.
+While empirically invariant to the choice of basis functions, its theoretical underpinnings remain open.
+A key limitation is reduced performance on highly volatile trajectories, where the objective is sensitive to rapid directional changes.
+Future work will focus on adaptive architectures and robust loss designs that explicitly capture signal complexity to enhance generalization.
 Appendix A Objective Function Gradient
@@
 $\begin{split}\left|\dot{\gamma}_{i}\left(\frac{k}{T}\right)-\dot{\hat{\gamma}}_{i}\left(\frac{k}{T}\right)\right|&\leq\left|\dot{\gamma}_{i}\left(\frac{k}{T}\right)-\frac{\gamma_{i}\left(\frac{k+1}{T}\right)-\gamma_{i}\left(\frac{k-1}{T}\right)}{\frac{2}{T}}\right|+\left|\frac{\gamma_{i}\left(\frac{k+1}{T}\right)-\gamma_{i}\left(\frac{k-1}{T}\right)}{\frac{2}{T}}-\frac{\hat{\gamma}_{i}\left(\frac{k+1}{T}\right)-\hat{\gamma}_{i}\left(\frac{k-1}{T}\right)}{\frac{2}{T}}\right|\\ &\leq\left|\dot{\gamma}_{i}\left(\frac{k}{T}\right)-\frac{\gamma_{i}\left(\frac{k+1}{T}\right)-\gamma_{i}\left(\frac{k-1}{T}\right)}{\frac{2}{T}}\right|+T\sup_{k}\left|\gamma_{i}\left(\frac{k}{T}\right)-\hat{\gamma}_{i}\left(\frac{k}{T}\right)\right|.\end{split}$
-For $10<\delta_{3}<$ 1, there exists a $T>0$ such that:
+For $0<\delta_{3}<1$, there exists a $T>0$ such that:
 $\left|\dot{\gamma}_{i}\left(\frac{k}{T}\right)-\frac{\gamma_{i}\left(\frac{k+1}{T}\right)-\gamma_{i}\left(\frac{k-1}{T}\right)}{\frac{2}{T}}\right|<\delta_{3},\quad\forall k.$
@@
 such that
-$\left|Q_{i}(\gamma_{i})\left(\frac{k}{T}\right)-Q_{i}(\hat{\gamma}_{i})\left(\frac{k}{T}\right)\right|\leq M_{-1}^{(3)}\varepsilon,\quad\text{where }M_{-1}^{(3)}\text{ is a constant.}$.
+$\left|Q_{i}(\gamma_{i})\left(\frac{k}{T}\right)-Q_{i}(\hat{\gamma}_{i})\left(\frac{k}{T}\right)\right|\leq M_{-1}^{(3)}\varepsilon,\quad\text{where }M_{-1}^{(3)}\text{ is a constant.}$
 This means at each fixed point $\frac{k}{T}$, the difference between $Q_{i}(\gamma_{i})$ and $Q_{i}(\hat{\gamma}_{i})$ is bounded by $M_{-1}^{(3)}\varepsilon$, and it is straightforward to have
@@
 These gradients can be expressed as products of gradients from ReLU, max pooling, etc., which are Lipschitz continuous.
+We note that both the 1D linear interpolation and a Fourier basis expansion in the stage of spectral representation in DeepFRC satisfy the Lipschitz continuity condition.
 Therefore, the theorem follows directly from Theorem 3.8 in Hardt et al.
@@
 Appendix E Experimental Details
-E.1 Description of Simulated Datasets
+E.1 Evaluation Metrics
+Registration Metrics:
+  • Registration Error ($\Delta Q_{\text{reg}}$): This metric, defined within the EFDA framework, quantifies the discrepancy between the estimated warping $\hat{\bm{\gamma}}$ and the ground-truth $\bm{\gamma}^{*}$.
+It is applicable only to simulated datasets where $\bm{\gamma}^{*}$ is known.
+Lower values indicate better alignment.
+[…]
+Reconstruction Metric:
+[…]
+Classification Metrics:
+  • Accuracy (ACC): The proportion of correctly classified samples.
+  • Macro-averaged $F_{1}$-score: Provides a balanced measure of performance across classes, especially important for imbalanced datasets [Sokolova and Lapalme, 2009].
+Higher values for both ACC and $F_{1}$-score indicate superior classification performance.
+E.2 Description of Simulated Datasets
 We construct a balanced two-class functional dataset $\{x_{i}(t),y_{i}\}_{i=1}^{N}$, where each function $x_{i}(t)$ for $t\in[0,1]$ is generated by composing amplitude and phase components.
@@
 where the parameters $\{a_{i}^{(cj)},\mu^{(cj)},\sigma_{i}^{(cj)}\}_{j=1}^{2}$ for $c=0,1$ are summarized in in the following table.
+We simulate $N=6000$ functions, each evaluated at 1000 time points.
 We split the balanced dataset into training, validation, and test sets with sizes 1600, 400, and 4000, respectively.
@@
 Index (01) (02) (11) (12)
-$a_{i}$ $13+U(-0.5,0.5)$ $12.5+U(-1,1)$ $12+U(-1,1)$ $13+U(-1.5,1.5)$
+$a_{i}$ 13+U(0.5,0.5)13+U(-0.5,0.5) 12.5+U(1,1)12.5+U(-1,1) 12+U(1,1)12+U(-1,1) 13+U(1.5,1.5)13+U(-1.5,1.5)
 $\mu$ 0.250 0.715 0.225 0.695
-$\sigma_{i}$ $0.06+U(-0.003,0.003)$ $0.075+U(-0.003,0.003)$ $0.06+U(-0.003,0.003)$ $0.1+U(-0.003,0.003)$
-[…]
-E.2 Description of Real Datasets
-The Four time series datasets from the UCR Time Series Classification Archive (https://www.timeseriesclassification.com/index.php):
+$\sigma_{i}$ 0.06+U(0.003,0.003)0.06+U(-0.003,0.003) 0.075+U(0.003,0.003)0.075+U(-0.003,0.003) 0.06+U(0.003,0.003)0.06+U(-0.003,0.003) 0.1+U(0.003,0.003)0.1+U(-0.003,0.003)
+Figure A1: Visualization of alignment by DeepFRC on simulated test data, with evaluation metrics $\rho=0.976$, $\Delta Q_{\text{reg}}=0.056$, ATV=0.504, $\text{ACC}=1.000$ and $F_{1}\text{-score}=1.000$.
+E.3 Description of Real Datasets
+The first four time series datasets are one-dimensional functional data, from the UCR Time Series Classification Archive (www.timeseriesclassification.com), and the last dataset is three-dimensional functional data, from Kaggle (www.kaggle.com):
 The Wave dataset [Liu and Yang, 2009] consists of eight simple gestures generated using accelerometers, collected through a particular procedure.
@@
 Each sample has 398 time points.
-E.3 Introduction of Baselines
+The MotionSense [Malekzadeh et al., 2019] dataset involves multivariate time-series signals recorded from smartphone sensors during six daily activities performed by 24 participants.
+From this dataset, we selected three signal channels (G.x, G.y and G.z) from two actions (walk and jog) to construct a binary classification task, containing 80 samples with 200 time points each.
+The train/validation/test split is 56/8/16.
+E.4 Introduction of Baselines
 Here we provide a brief introduction of the compared baseline methods:
@@
   • ADAFNN [Yao et al., 2021]: A model that inputs entire functional curves, adapting the bases during learning.
-E.4 Model Implementation Details on Real Data
+  • TSLANet [Eldele et al., 2024]: A model that inputs entire functional curves, with ASB (Adaptive Spectral Blocks) and ICB (Interactive Convolutional Blocks) as core structure.
+E.5 Model Implementation Details on Real Data
 For all real-world datasets, the functional input is Z-score standardized entry-wise based on the mean function and standard deviation.
@@
     Symbol (3 classes) 10 10 1e-3 1e-3
+    MotionSense 15 1 1e-3 1e-3
     In the discussion of stability with respect to basis expansion for DeepFRC, we compute Fourier, B-spline, and Polynomial scores as follows: The Fourier scores are obtained using the fast Fourier transform function rfft from torch.nn.fft, extracting the first 100 scores.
@@
 We use the default architecture for the models FuncNN and ADAFNN.
+  • The model TSLANet takes univariate or multivariate time series as input, first processing them via patch segmentation and positional embedding to retain temporal order.
+Its core structure consists of Adaptive Spectral Blocks (for dependency capture and denoising) and Interactive Convolutional Blocks.
+We use the default architecture for the model TSLANet.
+  • A note on handling multidimensional functional data is necessary.
+Models including DeepFRC, TTN, TSLANet, and FuncNN natively support such inputs.
+For other baselines, we implemented the following adaptations: SrvfRegNet was extended to process multidimensional data analogously to DeepFRC.
+For FCNNraw and ADAFNN, we concatenated signals from all channels into a single one-dimensional vector.
+For FCNNfourier, we used the concatenated basis expansion coefficients from each channel as the model input.
 Figure A2: Training loss vs. epochs by DeepFRC across the four real-life datasets.
 Figure A3: Visualization comparison of alignment performance by DeepFRC, TTN and SrvfRegNet across three two-class real datasets
-Figure A4: Performance of DeepFRC with different missing rates for raw data across four real-world datasets.
-Table A3: Robustness under data scarcity. $N$ and $n$ represent sample size and sequence length, respectively.
+Figure A4: Visualization comparison of alignment performance by DeepFRC, TTN and SrvfRegNet on MotionSense dataset in 3 channels G.x, G.y and G.z.
+Figure A5: Performance of DeepFRC with different missing rates for raw data across five real-world datasets.
+Table A3: Representation coefficients correlation on different noise level on synthetic data
+Noise level $\varepsilon\sim N(0,\sigma^{2})$
+Model $\sigma=0$ $\sigma=0.05$ $\sigma=0.10$ $\sigma=0.20$
+DeepFRC 0.983 0.982 0.979 0.969
+TTN 0.541 0.529 0.501 0.490
+SrvfRegNet 0.969 0.966 0.961 0.950
+Table A4: Robustness under data scarcity. NN and $n$ represent sample size and sequence length, respectively.
 Sparse samples ($n=1000$) Sparse time points ($N=6000$) Sparse samples and time points $(N,n)$
 Evaluation $N=200$ $N=100$ $N=50$ $n=200$ $n=100$ $n=50$ $(200,200)$ $(100,100)$ $(50,50)$
-ATV 0.605 2.105 2.571 0.580 0.784 0.867 1.097 2.454 12.492
-ACC 100.0% 99.3% 98.2% 100.0% 100.0% 100.0% 100.0% 95.0% 80.0%
-$\textrm{F}_{1}$-score 1.000 0.992 0.980 1.000 1.000 1.000 1.000 0.947 0.750
+ATV 0.580 0.784 0.867 0.605 2.105 2.571 1.097 2.454 12.492
+ACC 100.0% 100.0% 100.0% 100.0% 99.3% 98.2% 100.0% 95.0% 80.0%
+$\textrm{F}_{1}$-score 1.000 1.000 1.000 1.000 0.992 0.980 1.000 0.947 0.750